### v2

# Superfast Low Rank Approximation

Abstract
Low rank approximation of a matrix (hereafter LRA) is a highly important area of Numerical Linear and Multilinear Algebra and Data Mining and Analysis. One can operate with an LRA superfast – by using much fewer memory cells and flops than an input matrix $M$ has entries.111“Flop” stands for “floating point operation”. Can we, however, compute an LRA of a matrix superfast? Yes and no. For worst case inputs, any LRA algorithm fails miserably unless it involves all entries of $M$, but in computational practice worst case inputs seem to appear rarely, and accurate LRA are routinely obtained superfast for large and important classes of matrices. We test numerically a number of superfast algorithms for various synthetic and real world matrices and consistently arrive at quite accurate LRA. Adequate formal support for these empirical observations is still a challenge, but our upper estimates for the output accuracy of our superfast deterministic LRA algorithms show that they only fail for a very narrow input class and moreover that the spectral and Frobenius error norms of the outputs of our randomized LRA algorithms are expected to stay within a noticeable but reasonable factor from their optima.
Key Words:
Low-Rank Approximation (LRA), CUR LRA, Sublinear algorithms.
2020 Math. Subject Classification:
65Y20, 65F55, 68Q25, 68W20
1 Introduction
1.1. Superfast LRA CUR LRA: background. LRA of a matrix is among the most fundamental problems of Numerical Linear and Multilinear Algebra and Data Mining and Analysis, with applications ranging from machine learning theory and neural networks to term document data and DNA SNP data (see surveys [39, 44, 46, 69]). For example, matrices that represent Big Data (e.g., unfolding matrices of multidimensional tensors) tend to be so immense that only a tiny fraction of the entries fits primary memory of a computer, but quite typically they admit LRA, that is, can be closely approximated with the products of a few matrices of smaller sizes [72]. Such products can be multiplied by a vector superfast, that is, at sublinear cost – by using much fewer scalars and flops than the matrix has entries. [^2] Superfast algorithms have been studied extensively for Toeplitz, Hankel, Vandermonde, Cauchy, and other structured matrices having small displacement rank (see more in Sec. 1.5), and we extend this concept to computations of LRA of matrices that admit LRA. This implies dramatic acceleration of various popular matrix computations reduced to that basic operation, but can we compute an LRA of a matrix superfast or at least quasi superfast, that is, by involving only a small fraction of all entries of an input matrix?
Yes and No. No, because on the worst case inputs and even on the small matrix families of Example A.1 any algorithm fails miserably for computing LRA and even for a posteriori estimation of its accuracy unless it involves all entries of an input matrix. Yes, because in computational practice worst case inputs seem to appear rarely, and accurate LRA are routinely computed superfast for a large and important class of matrices. Most popular among these algorithms are deterministic Cross-Approximation (C–A) iterations, specializing Alternating Directions Implicit (ADI) Method [63] to LRA and recalled in Sec. 3.2. They output LRA in a memory efficient form of CUR, which preserves sparsity and non-negativity of an input matrix and is widely applied in data analysis.
Given an $m\times n$ matrix $M$ and a target rank 333“In practice, the target rank is rarely known in advance. $\dots$ LRA algorithms are usually implemented in an adaptive fashion $\dots$ until the error norm satisfies the desired tolerance” [70, Sec. 2.4]. $r$, one can fix a pair of small integers $k$ and $l$,
$r\leq k\leq m,~r\leq l\leq n,$ (1.1)
and then fix or randomly sample a pair of sets of $l$ columns and $k$ rows of $M$. The submatrices made up of these columns and rows serve as the factors $C$ and $R$ of a CUR of $M$, respectively, and share a $k\times l$ submatrix $G$ of $M$, said to be the generator of CUR. Given a generator $G$, one can compute a factor $U$ by using about $kl$ scalars and $O(kl\min\{k,l\})$ flops (see Sec. 2), that is, quasi superfast where $kl\ll mn$ and superfast where $kl\min\{k,l\}\ll mn$.
Our formal and empirical study of superfast LRA includes some novel algorithms and techniques. [^4] Our link of maxvol CUR rank-1 algorithm to matrix norm estimation in Sec. 3.3 can be of independent interest.
1.2. Our contribution I: the study of CUR LRA. In Part I we seek CUR LRA deterministically. In Table 5.1 we display the results of our numerical test of C-A iterations and some other superfast algorithms for computing CUR for various $n\times n$ synthetic and real world matrices for $n=256,512,1024$. We always had spectral output error norm within a factor of 30 from optimal with a single exception for the Primitive algorithm that uniformly at random sampled $k$ rows and $l$ columns of $M$ – to define factors $C$ and $R$ and a generator $G$ of CUR.555 Instead of sampling one can fix the rows and columns of $M$, e.g., the first ones, the last ones, or equally spaced ones among all $m$ rows and $n$ columns. Its output error norm has exceeded the optimal one by factors ranging from 1,000 to 40,000.
Adequate formal support for these empirical observations is a challenge, but we partly met in Thm. 4.1 (our Main Result I). According to this theorem any CUR LRA is within a factor of $(v+1)^{2}$ from optimal under the spectral norm $||\cdot||$ for
[…]
For comparison, Example A.1 shows that one cannot estimate MCUR||M-CUR|| superfast directly for a worst case matrix $M$ and a fixed $CUR$. Our upper bound of order of $(v+1)^{2}$ is too high to guide practical LRA but should be of qualitative interest. Our Tables 5.2 and 5.3 show empirically that superfast combination of the randomized near-optimal algorithms of [20] and C-A iterations tends to keep the output CUR LRA nearly as accurate as in the case of the randomized near-optimal algorithms of [20], running at superlinear cost.
1.3. Our contribution II: dual sketching algorithms. We provide another insight into the problem by deducing a little stronger error bound for a large class of superfast randomized LRA algorithms. With a high probability (whp) random sketching LRA algorithms [44, 39, 73, 70, 46, 69] compute near-optimal rank-$r$ approximation [^6] Here and hereafter “rank-$r$ approximation” means “approximation of rank at most rr”. of an $m\times n$ matrix $M$ by using auxiliary matrices (sketches) $FM$ and/or $MH$ for $k\times m$ and/or $n\times l$ random sketch matrices $F$ and/or $H$, respectively, for $r\leq\min\{k,l\}$, $\max\{k,l\}=r+p$, and an oversampling (overestimation) parameter $p>0$. For small positive integers $k$ and $l$ the algorithms run superfast, except for the computation of the sketches, and we seek superfast algorithms at that stage as well.
To simplify our presentation, we restrict it to real matrices throughout, with $M^{*}$ denoting transpose rather than Hermitian transpose of $M$, and to random sketching LRA that only uses sketches $MH$, as in [39], although in Sec. 7.2 we extend our results to algorithms using both sketches $FM$ and $MH$, and we can readily extend our study to complex matrices.
One can explore a variety of ways towards our goal. Maybe the simplest way is to randomly sample $l$ columns of the matrix $M$ or equivalently to let $H$ be the $n\times l$ leftmost submatrix of a random $n\times n$ permutation matrix under a fixed randomization model. Clearly, in this case only $nl$ elements of $M$ are accessed, and the calculation itself is given ”for free”. We can alternatively compute sketches $MH$ superfast by choosing Ultrasparse sketch matrices $H$. Example A.1 tells us that these ways (as well as any other superfast way) cannot yield accurate LRA for all inputs $M$, but in our numerical experiments with Ultrasparse sketch matrices we have quite consistently output accurate LRA for a large input class.
Towards explaining these test results and constructing a theory, we introduce a probabilistic structure in the space of input matrices $M$, thus defining dual sketching in comparison to customary primal one, where this matrix is fixed.
The output error norm of sketching algorithms of [39] is essentially defined by the product $V_{1}^{*}H$ where $V_{1}$ is the matrix of the $r$ top left singular vectors of $M$, associated with its $r$ top (largest) singular values. Now assume that $V_{1}^{*}$ is the Q factor in QR or QRP factorization of a Gaussian random matrix, filled with independent identically distributed normal (standard Gaussian) random variables, $H$ has orthonormal columns, and lrl-r is reasonably large. Then we readily prove that whp the sketching algorithm outputs LRA whose both Frobenius and spectral error norms are within a factor of $\sqrt{1+16~n/l}$ from optima; this is essentially Thm. 8.2 – our Main Result 2!
Thm. 8.2 can be applied to approximation by matrices of any fixed rank $r$ and cover random sketching algorithms that in the case of a proper Ultrasparse sketch matrix $H$ use $O((m+n)r^{2})$ flops, running superfast where $r^{2}\ll\max\{m,n\}$ and quasi superfast already where $r\ll\min\{m,n\}$.
1.4. Quality of dual LRA. Like Thm. 4.1, Thm. 8.2 should be of some qualitative interest, e.g., it implies that miserable failure of our superfast dual sketching algorithms is highly unlikely under our probabilistic model, although the algorithms fail for worst case inputs.
On the other hand, the upper bound of Thm. 8.2 is still too high, and we propose and successfully test some heuristic improvement of output accuracy (see Sec. 9):
(i) Thm. 8.2 holds for any matrix $H$ that has orthonormal columns or is just well-conditioned. To increase our chances for obtaining accurate LRA we apply sketching algorithms with a number of distinct sketch matrices $H$ or their combinations.
(ii) For sketch matrices obtained with sparse subspace embedding [9, 12, 14], [71, Sec. 3.3], [46, Sec. 9], one can devise LRA algorithms running at linear computational cost. According to [40], such acceleration versus LRA running at superlinear cost with Gaussian sketch matrices tends to make the output accuracy of the LRA somewhat less reliable, although [14] partly overcomes this problem for incoherent matrices, that is, filled with entries of comparable magnitude (see [14] or [11, Def. 1.1] for formal definition of incoherence). Multiplication by the matrix of Subsampled Randomized Hadamard or Fourier Transform (SRHT or SRFT) tends to make a matrix $M$ incoherent [14] but at superlinear cost. For a compromise, we devise Abridged SRHT sketch matrices. They are Ultrasparse and can be multiplied by a dense matrix superfast. Furthermore, they have orthonormal columns, and so Thm. 8.2 provides some probabilistic motivation for this otherwise heuristic LRA; our tests provide its empirical support. [^7] Abridged SRFT matrices can be defined similarly to Abridged SRHT matrices and empirically have similar power (see [55, Sec. 13.6 and Appendix B]); unlike real Abridged SRHT matrices they involve nonreal roots of unity.
1.5. Related work. For two decades of their worldwide application C-A (ADI-like) iterations have been consistently computing accurate CUR LRA superfast for a large and important matrix class (see [66, 36, 37, 34, 67, 2, 32, 6, 5, 30, 3, 33, 49, 1], and the bibliography therein). More recently, randomized superfast LRA algorithms have been proposed in [47, 8, 13] for some specified large and important classes of matrices defined by large numbers of independent parameters. Most notably, the authors of [47, 13] proved that their algorithms are expected to output near-optimal LRAs for Symmetric Positive Semidefinite (SPSD) matrices.
The superfast algorithms of [47, 8, 13] output LRA but not CUR LRA. The known algorithms extend any LRA to CUR LRA – at the price of increasing the spectral norm bounds on the output errors by a factor of $\sqrt{1+(n-r)r)}$ (see [39, Secs. 3.3.2 and 3.3.3]).
The sublinear cost of the algorithms of [47, 8, 13] does not include the superlinear cost of a posteriori estimation of their output error norms and correctness verification. The superfast deterministic algorithm of [42, Part III], however, computes CUR LRA of an $n\times n$ SPSD matrix with both spectral and Frobenius error norms within a factor of $n$ from optimal and as by-product, at no additional cost, estimates the output error norm, verifying correctness.
[43, Alg. 2.2] is a heuristic superfast CUR LRA algorithm, which is accurate for a large class of unsymmetric matrices according to the test results in [43]. The algorithm combines uniform choice of some subsets of the column and row sets for CUR with their updating by means of the Interpolative Decomposition of [39, Sec. 3.2.3] based on QR factorization with column pivoting. Jianlin Xia in [74] modified this algorithm by using Strong Rank Revealing (SRR) QR factorization of [26] instead of QR factorization with pivoting, called this technique “progressive alternating direction pivoting”, combined it with his additional techniques, in particular, randomized error estimation, and consistently obtained highly accurate LRAs for various real world inputs in his extensive numerical tests. Xia cited [42, 58] as his close predecessors, but our present algorithms rely on distinct techniques.
In formal study of CUR LRA, quite surprising was the theorem by Zamarashkin and Osinsky in [76] that for any $m\times n$ matrix $M$, any positive integer $k\leq\min\{m,n\}$, there exists a $k\times k$ CUR generator $G$ that defines CUR, which is optimal up to a factor of $k+1$ approximation of $M$ under the Frobenius matrix norm. In the next major step, Alice Cortinovis and Daniel Kressner [15] computed such a CUR at deterministic cost $O(kmn^{3})$ for $m\geq n$ by linking the problem to intricate subset selection. Superfast computation of a good generator, however, is a challenge, and the error norms of CUR LRAs of $M$ are unbounded over ill-conditioned generators $G$.
Chiu and Demanet in [11] proved that under the uniform sampling of row and column sets of $M$ the Primitive algorithm outputs accurate CUR LRA of $M$ whp, provided that there exists an accurate LRA $XYZ\approx M$, e.g., defined by the SVD of $M$, such that the matrices $X$ and $Z^{*}$ have orthonormal columns and are incoherent. [11] also extended these estimates to the case where only a single factor $X$ in some LRA $XYZ\approx M$ has orthonormal columns and is incoherent provided that the Primitive algorithm is extended with a single C-A iteration.
Alice Cortinovis and Lexing Ying [19] deduce the results of [11] under weaker assumptions about the matrices $X$ and $Z^{*}$ – they allow some of the column vectors to be sparse rather than incoherent if CUR LRA algorithm incorporates progressive alternating direction pivoting of [43, 74].
The highly important estimates of [11, 19] are still limited by the basic provision of incoherence of the factors of some unknown LRA of $M$. This provision does not hold for a large class of inputs and generally cannot be verified superfast for a fixed $M$ – in contrast to the bound of Thm. 4.1.
In Parts I and II we extend the unpublished preprints [54, 55, 56, 57], which cited [59, 60, 61] as their predecessors and stated motivation for further study of superfast LRA as their major goal; this is a major goal of our current work as well. Encouraged by the success of [74, 19] we revisit the two preprints [56, 57] and extract from them our Main Results I and II and the test results of Parts I and II, respectively.
1.6. Organization of our paper. In the next subsection we recall some definitions and auxiliary results. In Part I we cover computation of CUR LRA. We recall some background material in the next two sections. In Sec. 4 we estimate output errors of CUR LRA, proving our Main Result I. In Sec. 5 we cover numerical experiments for the algorithms of Part I. In Part II (Secs. 69) we study random sketching algorithms for LRA. In Secs. 6 and 7 we recall and slightly extend some known ones and recall their deterministic output error bounds, respectively. In Sec. 8 we prove error bounds for our dual LRA algorithms – our Main Result 2. In Sec. 9 we cover numerical tests of these algorithms. We devote Sec. 10 to conclusions. In Appendix A we specify some small families of matrices whose LRA fails unless it involves all entries of $M$. In Appendix B we first comment on some random sketching LRA and then define Abridged SRHT Ultrasparse matrices. In Appendix C we recall the known estimates for the norms of a Gaussian matrix and its pseudo inverse.
1.7. Some background for LRA. In this subsection, we write $|||\cdot|||$ to unify the spectral norm $||\cdot||$ and the Frobenius norm $||\cdot||_{F}$, by following [39, Thm. 9.1]. Hereafter $\mathbb{R}^{p\times q}$ denotes the class of $p\times q$ real matrices. For five positive integers $k,l,m,n,r$ satisfying Eqn. (1.1), the triplet of matrices
$X\in\mathbb{R}^{m\times k},~Y\in\mathbb{R}^{k\times\ell},~Z\in\mathbb{R}^{n\times\ell},~r\leq\min\{k,\ell\},$ (1.3)
as well as the product $XYZ^{*}$ represent a rank-$r$ approximation of a matrix $M\in\mathbb{C}^{m\times n}$ with the error matrix $E:=M-~XYZ^{*}$ (see Fig. 1).
[…]
Figure 1: LRA of $M$.
One can replace the triplet $\{X,Y,Z^{*}\}$ with the pair $\{XY,Z^{*}\}$ of $m\times\ell$ and $\ell\times n$ matrices or $\{X,YZ^{*}\}$ of $m\times k$ and $k\times n$ matrices.
An important 3-factor LRA of $M$ is its $r$-top SVD (see Fig. 2), $M_{r}=U_{r}\Sigma_{r}V_{r}^{*}$ for a diagonal matrix $\Sigma_{r}=\operatorname{diag}(\sigma_{j})_{j=1}^{r}$ of the $r$ largest singular values of $M$ and two matrices $U_{r}$ and $V_{r}$ of the $r$ associated left and right singular vectors, respectively. $M_{r}$ is said to be the $r$-truncation of $M$. It can be obtained from a matrix $M$ by setting to 0 all its singular values but the $r$ top (largest) ones. $r$-top SVD of $M$ is a compact SVD of $M_{r}$.
MM==00++$E$
Figure 2: The figure represents top SVD of a matrix.
Theorem 1.1.
(Eckart-Young-Mirsky Theorem, see [29, Thm. 2.4.8].) It holds that
[…]
under both spectral and Frobenius norms: $\tau_{r+1}(M)=\sigma_{r+1}(M)$ under the spectral norm, and $\tau_{r+1}(M)=\sigma_{F,r+1}(M):=\sqrt{\sum_{j>r}\sigma_{j}^{2}(M)}$ under the Frobenius norm.
Theorem 1.2.
[29, Cor. 8.6.2]. For a pair of $m\times n$} matrices $M$ and $M+E$ it holds that
[…]
$M^{+}$ denotes the Moore–Penrose pseudo inverse of $M$.
Lemma 1.1.
(The norm of the pseudo inverse of a matrix product, see, e.g., [34].) Suppose that $A\in\mathbb{R}^{k\times r}$, $B\in\mathbb{R}^{r\times l}$ and the matrices $A$ and $B$ have full rank $r\leq\min\{k,l\}$. Then
[…]
PART I: COMPUTATION OF CUR LRA
2 CUR decomposition and CUR LRA
For two sets $\mathcal{I}\subseteq\{1,\dots,m\}$ and $\mathcal{J}\subseteq\{1,\dots,n\}$, define the submatrices
$M_{\mathcal{I},:}:=(m_{i,j})_{i\in\mathcal{I};j=1,\dots,n},M_{:,\mathcal{J}}:=(m_{i,j})_{i=1,\dots,m;j\in\mathcal{J}},~{\rm and}~M_{\mathcal{I},\mathcal{J}}:=(m_{i,j})_{i\in\mathcal{I};j\in\mathcal{J}}.$
Given an $m\times n$ matrix $M$ of rank $r$ and its nonsingular $r\times r$ submatrix $G=M_{\mathcal{I},\mathcal{J}}$, one can readily verify that $M=M^{\prime}$ for
$M^{\prime}=CUR,~C=M_{:,\mathcal{J}},~U=G^{-1},~G=M_{\mathcal{I},\mathcal{J}},~{\rm and}~R=M_{\mathcal{I},:}.$ (2.1)
We call $G$ the generator and $U$ the nucleus of CUR decomposition of $M$ (see Fig. 1 for $X=C$, $Y=U$, $Z=R^{*}$, and $E=O$).
CUR decomposition is extended to CUR approximation of a matrix $M$, although the approximation $M^{\prime}\approx M$ for $M^{\prime}$ of (2.1) tends to be poor where the generator $G$ is ill-conditioned.
Remark 2.1.
The pioneering papers [36, 34, 37], as well as [32, 33, 30, 49], define CGR approximations having nuclei $G$; “G” can stand, say, for “germ”. We use the acronym CUR, which is more customary in the West. “U” can stand, say, for “unification factor”, but notice the alternatives of CNR, CCR, or CSR with $N$, $C$, and $S$ standing for “nucleus”, “core”, and “seed”.
By generalizing (2.1) we allow to use $k\times l$ CUR generators for $k$ and $l$ satisfying (1.1) and to choose any $l\times k$ nucleus $U$ for which the error matrix E=CURME=CUR-M has smaller norm.
Given two matrices $C$ and $R$, the minimal Frobenius error norm of CUR LRA
[…]
is reached for the nucleus $U=C^{+}MR^{+}$ (see [45, Eqn. (6)]). We, however, cannot compute such a nucleus superfast and instead seek canonical CUR LRA (cf. [20, 16, 49]) whose nucleus is the Moore-Penrose pseudo inverse of the $r$-truncation of a given CUR generator:
$U:=G_{r}^{+}.$ (2.2)
Given a generator $G$ we can compute $G_{r}^{+}$ by using about $kl$ scalars and $O(kl\min\{k,l\})$ flops.
2.1 CUR decomposition of a matrix
We do not use the following result, but it can be of independent interest.
Theorem 2.1.
[A necessary and sufficient criterion for CUR decomposition.] Let $M^{\prime}=CUR$ be a canonical CUR of $M$ for $U=G_{r}^{+}$, $G=M_{\mathcal{I},\mathcal{J}}$. Then $M^{\prime}=M$ if and only if $\operatorname{rank}(G)=\operatorname{rank}(M)$.
Proof.
$\sigma_{j}(G)\leq\sigma_{j}(M)$ for all $j$ because $G$ is a submatrix of $M$. Hence $\epsilon$-$(G)\leq\epsilon-rank$$(M)$ for all nonnegative $\epsilon$, and in particular $\operatorname{rank}(G)\leq\operatorname{rank}(M)$.
Now let $M=M^{\prime}=CUR$. Then clearly
$\operatorname{rank}(M)\leq\operatorname{rank}(U)=\operatorname{rank}(G_{r}^{+})=\operatorname{rank}(G_{r})\leq\operatorname{rank}(G),$
Hence
$\operatorname{rank}(G)\geq\operatorname{rank}(M),~{\rm and~so}~\operatorname{rank}(G)=\operatorname{rank}(M)~{\rm if}~M^{\prime}=M.$
It remains to deduce that
$M=CG_{r}^{+}R~{\rm if}~\operatorname{rank}(G)=\operatorname{rank}(M):=r,$
but in this case $G_{r}=G$, and so
$\operatorname{rank}(CG_{r}^{+}R)=\operatorname{rank}(C)=\operatorname{rank}(R)=r.$
Hence the rank-$r$ matrices $M$ and $CG_{r}^{+}R$ share their rank-$r$ submatrices $C$ and $R$. ∎
Remark 2.2.
Can we extend the theorem by proving that $M^{\prime}\approx M$ if and only if $\epsilon$-$(G)=\epsilon-rank$$(M)$ for a small positive $\epsilon$? The “only if” claim cannot be extended, e.g., for
$M=\begin{pmatrix}1&0&0\\ 0&\epsilon&0\\ 0&1&0\end{pmatrix}$,
$\epsilon\approx 0$, and the $2\times 2$ leading submatrix $G$ of $M$. Indeed, $\operatorname{rank}(M)=\operatorname{rank}(G)=2$, and so Thm. 2.1 implies that $M^{\prime}=M$, while $\epsilon$-$\operatorname{rank}(M)=2>\epsilon$-$\operatorname{rank}(G)=1.$
3 Superfast CUR LRA algorithms
Next we specify some superfast CUR LRA algorithms as heuristic but estimate their output accuracy in Thm. 4.1.
3.1 Primitive and Cynical algorithms
Given five integers $k$, $l$, $m$, $n$, and $r$ satisfying (1.1) and an $m\times n$ matrix $M$, sample uniformly or fix a pair of sets $\mathcal{I}$ and $\mathcal{J}$ of $k$ row and $l$ column indexes, respectively, to define a CUR generator $G$, compute its $r$-truncation $G_{r}$, build on it a canonical CUR LRA of $M$, and call this LRA algorithm Primitive. Apart from the selection of the sets $\mathcal{I}$ and $\mathcal{J}$ the algorithm amounts to computing the $r$-truncation $(M_{\mathcal{I},\mathcal{J}})_{r}$ of the matrix $M_{\mathcal{I},\mathcal{J}}$ and its pseudo inverse $((M_{\mathcal{I},\mathcal{J}})_{r})^{+}$; this involves about $kl$ scalars and $O(kl\min\{k,l\})$ flops.
The following CUR LRA algorithm (we call it Cynical) [^8] Here we allude to the benefits of the austerity and simplicity of primitive life, advocated by Diogenes the Cynic, and not to shamelessness and distrust associated with modern cynicism. first fixes or samples uniformly a larger submatrix $M^{\prime}$ of $M$ and then compresses it into a $k\times l$ CUR generator of $M$ by applying one of the algorithms of [26, 50, 20].
Figure 3: A cynical CUR algorithm (the stripes mark a $p\times q$ submatrix; a $k\times l$ CUR generator is shown in black).
Algorithm 3.1.
For an $m\times n$ matrix $M$, a target rank $r$, and four integers $k$, $l$, $p$, $q$ such that
$0<r\leq k\leq p\leq m,~r\leq l\leq q\leq n,~{\rm and}~kl<pq,$ (3.1)
fix or randomly sample a pair of sets $\mathcal{I}$ and $\mathcal{J}$ of $p$ row and $q$ column indexes, respectively, compute a $k\times l$ CUR generator $G_{k,l}$ for the $p\times q$ submatrix $M_{\mathcal{I},\mathcal{J}}$ by applying to it one of the algorithms of [26, 50, 20], compute the $r$-truncation $G=G_{k,l,r}$, and build on it CUR LRA of $M$.
The algorithm uses about $pq$ scalars and $O(pq\min\{p,q\})$ flops and turns into the Primitive algorithm for $p=k$ and $q=l$.
3.2 Cross-Approximation (C–A) iterations
C–A iterations (see Fig. 4):
  • For an $m\times n$ matrix $M$ and target rank $r$, fix four integers $k$, $l$, $p$ and $q$ satisfying (3.1). [C-A iterations are simplified in a special case where $p:=k$ and $q:=l$.]
  • Fix an $m\times q$ “vertical” submatrix of the matrix $M$, made up of its $q$ columns, [^9] One can alternatively begin C–A iterations with a “horizontal” submatrix. and apply to it one of the algorithms of [26, 50, 20], which output a $p\times q$ submatrix $G$.
  • Apply to it the Cynical algorithm and extend its output to CUR LRA of $M$.
  • Swap $p$ and $q$ and reapply the algorithm to the matrix $M^{*}$.
    [This is equivalent to computing a $k\times l$ CUR generator of a fixed $p\times n$ “horizontal” submatrix $M_{1}$ of $M$ that covers the submatrix $G$.]
  • Recursively alternate such “vertical” and “horizontal” steps until the new CUR generator computed at the current iteration coincides with the original one or until the number of recursive C–A steps exceeds a fixed tolerance bound. In both cases output the current CUR LRA of $M$.
Computational Cost. At its every “vertical” (resp. “horizontal”) step the algorithm involves about $mq$ (resp. $nl$) scalars and $O(mq^{2})$ (resp. $O(nl^{2})$) flops. For $p=q=r$ the algorithm of [30] uses $O(mr)$ (resp. $O(nr)$) flops per a “vertical” (resp. “horizontal”) step. Extensive effort has been invested in estimating the required number of C-A iterations (see [32, 30, 49, 1] and the references therein), but adequate support of their superfast empirical performance is still a challenge. [^10] Goreinov et al proved in [30] that every C-A step of maxvol (their variant of C-A iterations in the case where $p=q=r$) monotone increases the volume $|\det(G)|$, and reasonable bounds on the output error norms at a global maximum of the volume have been proved in [32, 49, 1]. Proving that maxvol yields global maximum superfast is still a challenge, however.
Figure 4: The first three recursive C–A steps output three striped submatrices.
Initialization recipes. Adaptive C-A iterations are initialized based on Gaussian elimination with complete pivoting (at a superlinear cost) combined with dynamic search for gaps in the spectrum of the singular values of $M$. Empirically the iterations converge very fast for a large and important class of input coming from the study of ODE and PDE (cf. [2, 3, 5, 6]). Maxvol variant of C-A iterations of [30] for $q=p=k=l=r$ is initialized by using $O(nr^{2})$ flops and still empirically converges fast for a large class of inputs, although the authors suggest using Gaussian elimination with complete pivoting if their initialization recipe does not work. One can try various other heuristic initialization recipes, e.g., based on progressive alternating direction pivoting of the papers [43, 74, 19].
3.3 Maxvol for rank-1 approximation and 1- and $\infty$-norm estimation
For a vector ${\bf v}=(v_{i})^{n}_{i=1}\in\mathbb{R}^{n}$ and a matrix $M=(m_{i,j})_{i,j=1}^{m,n}\in\mathbb{R}^{m\times n}$ write
[…]
Let $|M|:=\max_{i,j=1}^{m,n}|m_{i,j}|$ define the inconsistent matrix norm [^11] E.g., $|M^{2}|=n>|M|^{2}=1$ for $M=(1)_{i,j=1}^{n,n}$. $|M|:=\max_{i,j=1}^{m,n}|m_{i,j}|$, called Chebyshev’s norm in [49], such that
$|M|\leq||M||\leq\sqrt{mn}~|M|,~|M|\leq|||M||_{1}\leq m|M|,~|M|\leq||M||_{\infty}\leq n|M|.$ (3.3)
One can compute the norms $|M|$, $_{1}, M||M||_{\infty}$ by using 3mn33mn-3 additions and comparisons.
For $p=q=r=1$ maxvol computes an absolutely largest entry of $M=(m_{i,j})_{i,j=1}^{m,n}$ in its fixed column or row (in its “vertical” or “horizontal” step), respectively, by using m1m-1 or n1n-1 comparisons, respectively. Maxvol stops at the entry $m_{i,j}$ that maximizes both $|m_{g,j}|$ over all $g$ and $|m_{i,h}|$ over all $h$. For a large class of matrices $M$ this is also the global maximum $|m_{i,j}|=|M|$.
In that case Eqns. (3.2) and (3.3) imply that the 1-norms of the $j$th column vector $C:=(m_{g,j})_{g=1}^{m}$ and the $i$th row vector $R:=(m_{i,h})_{h=1}^{n}$ estimate the 1-norm and the $\infty$-norm of $M$, respectively, being off by at most factors of $m$ and $n$, respectively.
Furthermore, write $U:=1/m_{i,j}$, obtain rank-1 approximation of $M$ with the matrix $CUR$, and recall from the result of [1], extending [32, 49], that
$|M-CUR|\leq\frac{2\sigma_{1}(M)\sigma_{2}(M)}{(\sigma_{1}^{2}(M)+\sigma^{2}_{2}(M))^{1/2}}.$
Thm. 1.1 and bound (3.3) combined imply that this approximation is optimal under the spectral norm up to a factor of $2\sqrt{mn}$.
4 A posteriori errors of a canonical CUR LRA
Example A.1 shows that a posteriori errors of an LRA given by a matrix $M^{\prime}\approx M$ cannot be estimated superfast for worst case inputs, but next we prove such a bound where $M^{\prime}$ is given by a triplet $\{C,U,R\}$.
4.1 Error Estimation: an Outline and the Statement
We estimate the error norm of CUR rank-$r$ approximation of $M$ by comparing it with CUR decomposition of a nearby rank-$r$ matrix $M^{\prime}$ defined by the same sets of row and column indices.
Outline 4.1.
[Error Estimation for a Canonical CUR LRA.]
  1. 1.
    Consider (but do not compute) an auxiliary $m\times n$ matrix $M^{\prime}$ of rank $r$ that approximates the matrix $M$ within a fixed norm bound $\epsilon$ such that
[…]
    [E.g., we can choose $M^{\prime}=M_{r}$ and $\epsilon:=\sigma_{r+1}(M)$.]
  2. 2.
    For the matrices $M$ and $M^{\prime}$ fix two row and column index sets $\mathcal{I}$ and $\mathcal{J}$, respectively, and define $k\times l$ generators $G=M_{\mathcal{I},\mathcal{J}}$ and $G^{\prime}=M_{\mathcal{I},\mathcal{J}}^{\prime}$, nuclei $U=G_{r}^{+}$ and $U^{\prime}=G_{r}^{\prime+}$, and canonical CUR approximation $M\approx CUR$ and decomposition $M^{\prime}=C^{\prime}U^{\prime}R^{\prime}$.
  3. 3.
    Observe that
[…]
  4. 4.
    Bound the norm $C^{\prime}U^{\prime}R^{\prime}-CUR||$ in terms of the values $\epsilon$, $C||$, $U||$, and $R||$.
Next we elaborate upon step 4 provided that we have already performed steps 1 – 3.
Theorem 4.1.
Given an $m\times n$ matrix $M$, a $k\times l$ matrix $G:=M_{\mathcal{I},\mathcal{J}}$, a positive $\epsilon\leq\sigma_{r}(G)$, a positive integer $r<\min\{k,l\}$, and an $m\times n$ rank-$r$ matrix $M^{\prime}$ satisfying (4.1), fix $v$ of (1.2), write
$C:=M_{:,\mathcal{J}},~R:=M_{\mathcal{I},:},~U=G_{r}^{+},$ (4.3)
$\zeta:=\begin{cases}\sqrt{2}&{\rm for}~~r=\min\{k,l\}\\ (1+\sqrt{5})/2&{\rm for}~~r<\min\{k,l\}\end{cases}$ (4.4)
[…]
$M-CUR||\leq(v+1)\Big{(}\frac{2\zeta}{1-\theta}~(v+1)+2\Big{)}\epsilon,~{\rm where}~2\zeta\leq 1+\sqrt{5}.$ (4.5)
In view of (4.2) we only need to estimate the norm $C^{\prime}U^{\prime}R^{\prime}-CUR||$.
4.2 The first bound on $C^{\prime}U^{\prime}R^{\prime}-CUR||$
Lemma 4.1.
Fix five integers $k$, $l$, $m$, $n$, and $r$ such that $r\leq k\leq m$ and $r\leq l\leq n$, an $m\times n$ matrix $M$, its rank-$r$ approximation $M^{\prime}$ satisfying (4.1), and canonical CUR LRAs
$M\approx CUR~{\rm and}~M^{\prime}=C^{\prime}U^{\prime}R^{\prime}$
defined by the same pair of index sets $\mathcal{I}$ and $\mathcal{J}$ of cardinality $k$ and $l$, respectively, such that
$C:=M_{:,\mathcal{J}},~R:=M_{\mathcal{I},:},~U=G_{r}^{+},~~C^{\prime}:=M^{\prime}_{:,\mathcal{J}},~R^{\prime}:=M^{\prime}_{\mathcal{I},:},~U^{\prime}=G_{r}^{\prime+},$
$G=M_{\mathcal{I},\mathcal{J}},~{\rm and}~G^{\prime}=M^{\prime}_{\mathcal{I},\mathcal{J}}.$
Then
[…]
Proof.
Notice that
$CUR-C^{\prime}U^{\prime}R^{\prime}=(C-C^{\prime})UR+C^{\prime}U(R-R^{\prime})+C^{\prime}(U-U^{\prime})R^{\prime}.$
Therefore
[…]
4.3 Estimation of the norm $U-U^{\prime}||$
Lemma 4.2.
Under the assumptions of Lemma 4.1 we have
[…]
Proof.
Recall that $\operatorname{rank}(G^{\prime})=\operatorname{rank}(M^{\prime})=\operatorname{rank}(G_{r})=r$ and that
[…]
Then apply [4, Thm. 2.2.5] for $A=G,~B=G^{\prime}$. ∎
Lemma 4.3.
[…]
Proof.
Thm. 1.2 implies that
[…]
Substitute (1.2) and obtain that $\sigma_{r}(G^{\prime})\geq(1-\theta)\sigma_{r}(G)>0$. Hence
[…]
Corollary 4.1.
[…]
4.4 Proof of Thm. 4.1
Combine bound (4.2) and Lemma 4.1 and obtain
[…]
Combine (1.2) and Cor. 4.1 and deduce that
$||C^{\prime}||~||U-U^{\prime}||\leq(||C||+\epsilon)\frac{2\zeta}{1-\theta}~||U||^{2}\epsilon<\frac{2\zeta}{1-\theta}(||C||~|U||^{2}+||U||)\epsilon.$
Hence
[…]
Substitute relationships (4.3) and obtain
[…]
Combine this bound with Eqns. (4.6) and (4.7) and obtain the theorem.
5 Numerical experiments
5.1 Test overview
For both synthetic and real world input matrices the test results in our Table 5.1 show very good output accuracy of CUR LRA output by just five C-A loops (with uniform initialization), which was still quite good even where we applied the Cynical algorithm. The accuracy was significantly worse in the case of the Primitive algorithm. Tables 5.2 and 5.3 show the other test results for CUR LRA described at the end of Sec. 1.2.
We have performed the tests in the Graduate Center of the City University of New York by using MATLAB. Generating synthetic inputs we used Gaussian random matrices, filled with independent identically distributed normal (standard Gaussian) random variables. In particular we applied its standard normal distribution function “randn()” to generate Gaussian matrices and calculated $\epsilon$-ranks of matrices for $\epsilon=10^{-6}$ by using the MATLAB’s function ”rank(-,1e-6)”, which only counts singular values greater than $10^{-6}$. Some numerical experiments were executed with software custom programmed in $C^{++}$ and compiled with LAPACK version 3.6.0 libraries.
Our tables display the mean value of the relative output errors $\frac{\|M-\tilde{M}\|}{\|M\|}$ for an input matrix $M$ and output $\tilde{M}:=CUR$ and the standard deviation (std) over 1000 runs for every class of inputs.
5.2 Input matrices for LRA
We used the following two classes of input matrices $M$ for testing LRA algorithms.
Class I (Synthetic inputs): $n\times n$ random matrices $M$ having expected rank $r$, that is,
$M=G_{1}*G_{2}+10^{-10}G_{3},$
for three Gaussian matrices $G_{1}$ of size $n\times r$, $G_{2}$ of size $r\times n$, and $G_{3}$ of size $n\times n$.
[…]
namely, to the following six input families from the Database:
baart: Fredholm Integral Equation of the first kind,
shaw: one-dimensional image restoration model,
gravity: 1-D gravity surveying model problem,
wing: problem with a discontinuous solution,
foxgood: severely ill-posed problem,
inverse Laplace: inverse Laplace transformation.
5.3 Four test algorithms
In our tests we applied and compared the following four algorithms for computing CUR LRA to input matrices $M$ having $\epsilon$-rank $r$:
  • Tests 1 (The Primitive algorithm for $k=l=r$): Uniformly choose two index sets $\mathcal{I}$ and $\mathcal{J}$, both of cardinality $r$, then compute a nucleus $U=M_{\mathcal{I},\mathcal{J}}^{-1}$ and define CUR LRA
    $\tilde{M}:=CUR=M_{:,\mathcal{J}}\cdot M_{\mathcal{I},\mathcal{J}}^{-1}\cdot M_{\mathcal{I},\cdot}.$ (5.1)
  • Tests 2 (Five loops of C–A): Uniformly choose an initial row index set $\mathcal{I}_{0}$ of cardinality $r$, then perform five loops of C–A (cf. Sec. 3.2) incorporating Alg. 1 of [50] as a subalgorithm that produces $r\times r$ CUR generators. At the end compute a nucleus $U$ and define CUR LRA as in Tests 1.
  • Tests 3 (A Cynical algorithm for $p=q=4r$ and $k=l=r$): Uniformly choose a row index set $\mathcal{K}$ and a column index set $\mathcal{L}$, both of cardinality $4r$, and then apply Algs. 1 and 2 from [50] to compute an $r\times r$ submatrix $M_{\mathcal{I},\mathcal{J}}$ of $M_{\mathcal{K},\mathcal{L}}$. Compute a nucleus and obtain CUR LRA by applying equation (5.1).
  • Tests 4 (Combination of a single C–A loop with Tests 3): Uniformly choose a column index set $\mathcal{L}$ of cardinality $4r$; then perform a single C–A loop (made up of a single horizontal step and a single vertical step): First by applying Alg. 1 from [50] define an index set $\mathcal{K}^{\prime}$ of cardinality $4r$ and the submatrix $M_{\mathcal{K}^{\prime},\mathcal{L}}$ in $M_{:,\mathcal{L}}$; then by applying this algorithm to matrix $M_{\mathcal{K}^{\prime},:}$ find an index set $\mathcal{L}^{\prime}$ of cardinality $4r$ and define submatrix $M_{\mathcal{K}^{\prime},\mathcal{L}^{\prime}}$ in $M_{\mathcal{K}^{\prime},:}$. Then proceed as in Tests 3 – find an $r\times r$ submatrix $M_{\mathcal{I},\mathcal{J}}$ in $M_{\mathcal{K}^{\prime},\mathcal{L}^{\prime}}$ by applying Algs. 1 and 2 from [50], compute a nucleus and CUR LRA.
5.4 CUR LRA of the matrices of class I
Table 5.1 shows the summary statistics for each of the four tests matrices of class I for $n=256,512,1024$ and $r=8,16,32$. In the column SVD we display the data for the best rank-$r$ approximation given by the $(r+1)$-st largest singular value to establish the baseline for the performance of our tests. The results of Tests 1 fall in the range $10^{-8},10^{-7}]$, those of Tests 2, 3 and 4 in $10^{-11},10^{-10}]$ in the decreasing order of accuracy. On the average the error bounds of Tests 2, 4, 3, and 1 exceed the baseline bounds in the column SVD by roughly factors of 10, 15, 30, and over 10,000, respectively, but even the crudest LRAs output in Test 1 could be used in some applications.
SVD Tests 1 Tests 2 Tests 3 Tests 4
n r mean std mean std mean std mean std mean std
256 8 1.01e-11 3.92e-13 1.60e-08 8.65e-08 5.94e-11 8.06e-12 1.13e-10 2.36e-11 8.23e-11 1.64e-11
256 16 9.12e-12 3.09e-13 2.44e-07 5.85e-06 7.31e-11 1.02e-11 1.12e-10 2.02e-11 9.45e-11 1.51e-11
256 32 7.80e-12 2.32e-13 4.82e-08 2.50e-07 8.93e-11 1.10e-11 1.13e-10 1.69e-11 1.04e-10 1.50e-11
512 8 7.64e-12 2.23e-13 3.50e-08 2.72e-07 5.71e-11 7.08e-12 1.21e-10 2.67e-11 8.34e-11 1.62e-11
512 16 7.06e-12 1.68e-13 1.18e-07 2.53e-06 7.08e-11 8.96e-12 1.26e-10 2.21e-11 9.98e-11 1.59e-11
512 32 6.36e-12 1.37e-13 7.43e-08 7.47e-07 9.25e-11 1.14e-11 1.34e-10 1.92e-11 1.20e-10 1.73e-11
1024 8 5.63e-12 1.13e-13 2.42e-08 2.46e-07 5.39e-11 5.89e-12 1.28e-10 2.83e-11 8.10e-11 1.55e-11
1024 16 5.34e-12 9.23e-14 6.12e-08 6.91e-07 6.94e-11 7.68e-12 1.37e-10 2.35e-11 1.04e-10 1.73e-11
1024 32 4.95e-12 7.55e-14 6.20e-07 1.36e-05 9.17e-11 1.06e-11 1.51e-10 2.09e-11 1.29e-10 1.87e-11
Table 5.1: Errors of CUR LRA of random matrices of class I
5.5 LRA by means of random sampling and C-A acceleration
Whp the randomized algorithms of [20] run at superlinear cost and are expected to output an LRA within any fixed positive relative error norm bound $\epsilon$. We tested both the randomized algorithm of [20] and its combination with C-A iterations. Tables 5.2 and 5.3 display the relative errors $\frac{\|M-\tilde{M}\|}{\|M\|}$ of the LRA $\tilde{M}$ of the matrices $M$ computed in two ways: by means of [20, Alg. 2] (see the lines marked “CUR”) or eight C-A iterations with [20, Alg. 1] applied at all vertical and horizontal steps (see the lines marked “C-A”). The overall cost of performing the algorithms is superlinear in the former case and sublinear in the latter case. The columns of the tables marked with $\epsilon$-rank” display $\epsilon$-rank of an input matrix. The columns of the tables marked with ”k=lk=l” show the number of rows and columns in a square matrix of CUR generator.
As we pointed out at the end of Sec. 1.2, in almost all our tests this dramatic acceleration was achieved at the price of only minor deterioration of output accuracy.
input algorithm m n $\epsilon$-rank k=l mean std
finite diff C-A 608 1200 94 376 6.74e-05 2.16e-05
finite diff CUR 608 1200 94 376 6.68e-05 2.27e-05
finite diff C-A 608 1200 94 188 1.42e-02 6.03e-02
finite diff CUR 608 1200 94 188 1.95e-03 5.07e-03
baart C-A 1000 1000 6 24 2.17e-03 6.46e-04
baart CUR 1000 1000 6 24 1.98e-03 5.88e-04
baart C-A 1000 1000 6 12 2.05e-03 1.71e-03
baart CUR 1000 1000 6 12 1.26e-03 8.31e-04
baart C-A 1000 1000 6 6 6.69e-05 2.72e-04
baart CUR 1000 1000 6 6 9.33e-06 1.85e-05
shaw C-A 1000 1000 12 48 7.16e-05 5.42e-05
shaw CUR 1000 1000 12 48 5.73e-05 2.09e-05
shaw C-A 1000 1000 12 24 6.11e-04 7.29e-04
shaw CUR 1000 1000 12 24 2.62e-04 3.21e-04
shaw C-A 1000 1000 12 12 6.13e-03 3.72e-02
shaw CUR 1000 1000 12 12 2.22e-04 3.96e-04
Table 5.2: LRA errors of Cross-Approximation (C-A) incorporating [20, Algorithm 1] in comparison to stand-alone CUR of [20, Algorithm 2].
input algorithm m = n $\epsilon$-rank $k=l$ mean std
foxgood C-A 1000 10 40 3.05e-04 2.21e-04
foxgood CUR 1000 10 40 2.39e-04 1.92e-04
foxgood C-A 1000 10 20 1.11e-02 4.28e-02
foxgood CUR 1000 10 20 1.87e-04 4.62e-04
wing C-A 1000 4 16 3.51e-04 7.76e-04
wing CUR 1000 4 16 2.47e-04 6.12e-04
wing C-A 1000 4 8 8.17e-04 1.82e-03
wing CUR 1000 4 8 2.43e-04 6.94e-04
wing C-A 1000 4 4 5.81e-05 1.28e-04
wing CUR 1000 4 4 1.48e-05 1.40e-05
gravity C-A 1000 25 100 1.14e-04 3.68e-05
gravity CUR 1000 25 100 1.41e-04 4.07e-05
gravity C-A 1000 25 50 7.86e-04 4.97e-03
gravity CUR 1000 25 50 2.22e-04 1.28e-04
inverse Laplace C-A 1000 25 100 4.15e-04 1.91e-03
inverse Laplace CUR 1000 25 100 5.54e-05 2.68e-05
inverse Laplace C-A 1000 25 50 3.67e-01 2.67e+$0$
inverse Laplace CUR 1000 25 50 2.35e-02 1.71e-01
Table 5.3: LRA errors of Cross-Approximation (C-A) incorporating [20, Algorithm 1] in comparison to stand-alone CUR of [20, Algorithm 2].
PART II: LRA VIA RANDOM SKETCHING
6 LRA by means of sketching: four algorithms
Next we slightly generalize the sketching LRA algorithms of [39, 70] (see Remarks 6.16.3).
Algorithm 6.1.
Input:
An $m\times n$ matrix $M$ and a target rank $r$.
Output:
Two matrices $X\in\mathbb{R}^{m\times l}$ and $Y\in\mathbb{R}^{l\times n}$ for $r\leq l\leq n$ defining an LRA $\tilde{M}=XY$ of $M$.
Initialization:
Fix an integer $l=r+p\leq n$, for $p\geq 0$, and an $n\times l$ matrix $H$ of full rank $l$.
Computations:
  1. 1.
    Compute the $m\times l$ matrix $MH$.
  2. 2.
    Fix a nonsingular $l\times l$ matrix $T^{-1}$ and output the $m\times l$ matrix $X:=MHT^{-1}$.
  3. 3.
[…]
Algorithm 6.2.
Input:
As in Alg. 6.1.
Output:
Two matrices $X\in\mathbb{R}^{k\times n}$ and $Y\in\mathbb{R}^{m\times k}$ defining an LRA $\tilde{M}=YX$ of $M$.
Initialization:
Fix an integer $k=r+p\leq m$, for $p\geq 0,$ and a $k\times m$ matrix $F$ of full numerical rank $k$.
Computations:
  1. 1.
    Compute the $k\times m$ matrix $FM$.
  2. 2.
    Fix a nonsingular $k\times k$ matrix $S^{-1}$; then output $k\times n$ matrix $X:=S^{-1}FM$.
  3. 3.
[…]
The following algorithm combines row and column sketching.
Algorithm 6.3.
Input:
As in Alg. 6.1.
Output:
Two matrices $X\in\mathbb{R}^{m\times k}$ and $Y\in\mathbb{R}^{k\times n}$ defining an LRA $\tilde{M}=XY$ of $M$.
Initialization:
Fix two integers $k$ and $l$, $r\leq k\leq m$ and $r\leq l\leq n$; fix two matrices $F\in\mathbb{R}^{k\times m}$ and $H\in\mathbb{R}^{n\times l}$ of full numerical ranks and two nonsingular matrices $S\in\mathbb{R}^{k\times k}$ and $T\in\mathbb{R}^{l\times l}$.
Computations:
1. Output the matrix $X=MHT^{-1}\in\mathbb{R}^{m\times l}$.
2. Compute the matrices $U:=S^{-1}FM\in\mathbb{R}^{k\times n}$ and $W:=S^{-1}FX\in\mathbb{R}^{m\times l}$.
[…]
Remark 6.1.
We can obtain Algs. 6.2 by applying Alg. 6.1 to the transpose $M^{*}$. Likewise, by applying Alg. 6.3 to $M^{*}$ we obtain Alg. 6.4. We only study Algs. 6.1 and 6.3, but can readily extend that study to Algs. 6.2 and 6.4.
Remark 6.2.
Fix $k=r$, $l>k$, a random $n\times l$ matrix $H$ (e.g., Gaussian or SRHT matrix), identity matrix $S$, and $T$ equal to the $R$ factor in the $QR$ factorization of $MH$. Then the matrix $X$ has orthonormal columns and Algs. 6.1 turns into Proto-algorithm of [39, Sec. 9], while Alg. 6.3 turns into the Generalized Nyström algorithm of [48, Eqn. (3)]. [48] stabilizes the latter algorithm numerically – essentially by means of setting to 0 all singular values of the matrix $W$ exceeded by a fixed $\epsilon$, “a modest multiple of the unit roundoff $u$ times W||W||”. Our study can be readily extended to such a stabilized LRA because the stabilization little affects complexity of LRA and only improves its output accuracy. We can obtain various other modifications of Algs. 6.1 and 6.3 by fixing other sketching matrices $F$ and $H$.
Remark 6.3.
Column (resp. row) sketching turns into column (resp. row) subset selection where $H$ in Alg. 6.1 (resp. $F$ in Alg. 6.2) is a sampling matrix, that is, a full rank submatrix of a permutation matrix.
7 Deterministic output error bounds for sketching algorithms
7.1 Deterministic error bounds of Range Finder
Next we recall some known estimates for the errors of Alg. 6.1, to be used in the next section.
Theorem 7.1.
[39, Thm. 9.1]. Suppose that Alg. 6.1 has been applied to a matrix $M$ and let
$M=\begin{pmatrix}U_{1}&U_{2}\end{pmatrix}\begin{pmatrix}\Sigma_{1}&\\ &\Sigma_{2}\end{pmatrix}\begin{pmatrix}~V_{1}^{*}\\ ~V_{2}^{*}\end{pmatrix}~~{\rm and}~M_{r}=U_{1}\Sigma_{1}V_{1}^{*}$
be SVDs of the matrices $M$ and its rank-$r$ truncation $M_{r}$, respectively. $\Sigma_{2}=O$ and $XY=M$ if $\operatorname{rank}(M)=r$. The $r$ columns of $V_{1}$ are the $r$ top right singular vectors of $M$.] Write
$C_{1}=V^{*}_{1}H,~C_{2}=V^{*}_{2}H.$ (7.1)
Assume that $H||_{2}\leq 1$ and $\operatorname{rank}(C_{1})=r$. Then
[…]
Corollary 7.1.
Under the assumptions of Thm. 7.1 and for $\tau_{r+1}(M)$ of Thm. 1.1 it holds that
[…]
Proof.
The corollary follows from (7.2) because
[…]
7.2 Impact of pre-multiplication on the errors of LRA
The following theorem shows that the overall error bounds of Alg. 6.3 are dominated by the product of the norm $|W|||$ and the error norm bound of Alg. 6.1.
Theorem 7.2.
See [70]. Let Alg. 6.3 output a matrix $XY$ for $Y=(FX)^{+}FM$ and let $m\geq k\geq l=\operatorname{rank}(X)$. Then
$M-XY=W(M-XX^{+}M)~{\rm for}~W=I_{m}-X(FX)^{+}F,$ (7.4)
[…]
Proof.
Recall that $Y=(FX)^{+}FM$ and notice that $(FX)^{+}FX=I_{l}$ if $k\geq l=\operatorname{rank}(FX)$. Therefore, $Y=X^{+}M+(FX)^{+}F(M-XX^{+}M)$. Consequently, (7.4) and (7.5) hold. ∎
Remark 7.1.
[…]
if both matrices $F$ and $X$ have orthonormal columns. The sketch matrix $F$ is our choice, and for $X$ we ensure column orthogonality by properly choosing the matrix $T$ in Alg. 6.3 (see Remark 6.2).
8 Output error norm bounds for dual sketching algorithms
Given the matrices $MHT^{-1}$ and $S^{-1}FM$, Alg. 6.3 uses $O(kln)$ flops and hence runs superfast for $kl\ll m$. If also $l^{2}\ll m$ and $k^{2}\ll n$, then for proper Ultrasparse matrices $F$ and $H$ we can compute the matrices $MHT^{-1}$ and $S^{-1}FM$ superfast as well and hence can perform entire Algorithm 6.3 superfast, involving much less than $mn$ entries of $M$ and other scalars already where $k\ll m$ and $l\ll n$. Although such an LRA fails for a worst case input (cf. Appendix A), it succeeds whp in the case of any fixed, possibly sparse, well-conditioned matrix $H$ of full rank and a random input matrix $M$ that admits LRA, as we prove next. In view of the previous subsection we only need to prove this for Alg. 6.1.
8.1 Auxiliary results
Lemma 8.1.
Suppose that $G$ and $H$ are $r\times n$ and $n\times l$ matrices, respectively, $r<l<n$, GHGH has full rank $r$, and $Q$ is an $r\times n$ matrix with orthonormal rows such that $Q$ and $G$ have the same row space. Then
$\sigma_{r}(QH)\geq\frac{\sigma_{r}(GH)}{\sigma_{1}(G)}.$
Proof.
Without loss of generality, suppose that $R\in\mathbb{R}^{r\times r}$ and $G=RQ$. Then
$\sigma_{r}(QH)=\sigma_{r}(R^{-1}GH)\geq\sigma_{r}(GH)\cdot\sigma_{r}(R^{-1}).$
Hence $\sigma_{r}(R^{-1})=1/\sigma_{1}(G)$ because the matrices $R$ and $G$ share their singular values. ∎
Constant matrices are filled with constants, unlike random matrices, filled with random variables.
Theorem 8.1.
[Non-degeneration of a Gaussian matrix.] Suppose that $M\in\mathbb{R}^{p\times q}$ is a constant matrix, $r\leq\operatorname{rank}(M)$, and $F$ and $H$ are $r\times p$ and $q\times r$ independent Gaussian matrices, respectively. Then the matrices $F$, $H$, FMFM, and $MH$ have full rank $r$ with probability 1.
Proof.
Rank deficiency of matrices $H$, FMFM, and $MH$ is equivalent to turning into 0 the determinants $\det(FF^{*})$, $\det(H^{*}H)$, $\det((MH^{*})MH)$, and $\det(FM(FM)^{*})$, respectively. The claim follows because these equations define algebraic varieties of lower dimension in the linear spaces of the entries, considered independent variables (cf., e.g., [7, Prop. 1]). ∎
Remark 8.1.
Events that occur with probability 0 are immaterial for our probability estimates, and hereafter we say that a matrix has full rank even if it is rank deficient with probability 0.
Lemma 8.2.
[Orthogonal Invariance.] [68, Theorem 3.2.1]. Suppose that $G$ is an $m\times n$ Gaussian matrix, $k\leq\min\{m,n\}$ is a positive integer, and $S\in\mathbb{R}^{k\times m}$ and $T\in\mathbb{R}^{n\times k}$ are constant matrices, having orthonormal rows and columns, respectively. Then $SG$ and $GT$ are random matrices having distribution of $k\times n$ and $m\times k$ Gaussian random matrices, respectively.
Hereafter write $e:=2.7182818\dots$.
Lemma 8.3.
Suppose that $G$ is an $r\times n$ Gaussian matrix, $H$ is an $n\times l$ matrix with orthonormal columns, $n>l>r$, $l\geq 4$, $Q$ is a matrix with orthonormal rows, and $Q$ and $G$ share their row space. Fix two positive parameters $t_{1}$ and $t_{2}<1$. Then
$\sigma_{r}(QH)\geq t_{2}\cdot\frac{\sqrt{l}-\sqrt{r/l}+\sqrt{1/l}}{t_{1}+\sqrt{r}+\sqrt{n}}\cdot\frac{1}{e}\$
with a probability no less than $1-\exp(-t_{1}^{2}/2)-(t_{2})^{l-r}$.
Proof.
The matrix $GH$ has distribution of an $r\times l$ Gaussian random matrix by virtue of Lemma 8.2, and hence we can assume that it has full rank (see Remark 8.1). Now recall Thm. C.1 and claim (ii) of Thm. C.2 and obtain
$\operatorname{Probability}\{\sigma_{1}(G)>t_{1}+\sqrt{r}+\sqrt{n}\}<\exp(-t_{1}^{2}/2)~{\rm for}~t_{1}\geq 0~{\rm and}$
$\operatorname{Probability}\Big{\{}\sigma_{r}(GH)\leq t_{2}\cdot\frac{l-r+1}{e\sqrt{l}}\Big{\}}\leq(t_{2})^{l-r}~{\rm for}~t_{2}<1~{\rm and}~l\geq 4.$
Combine the latter two inequalities, the union bound, and Lemma 8.1 and obtain Lemma 8.3. ∎
Lemma 8.3 implies that $\sigma_{r}(QH)$ has at least order of $\sqrt{l/n}$ whp.
Corollary 8.1.
For $n,l,r,G,Q$, and $H$ of Lemma 8.3, let $n>36r$ and l>22(r1)l>22(r-1). Then
$\operatorname{Probability}\big{\{}\sigma_{r}(QH)\geq\frac{1}{4}\sqrt{l/n}\big{\}}>1-\exp\Big{(}-\frac{n}{72}\Big{)}-\exp\Big{(}-\frac{l-r}{20}\Big{)}.$
Proof.
Write $t_{1}:=\frac{1}{3}\sqrt{n}-\sqrt{r}$ and $t_{2}:=\frac{1}{3}~\frac{le}{l-r+1}$, recall that $n>36r$ and l>22(r1)l>22(r-1), and then readily verify that $t_{1}>\frac{\sqrt{n}}{6}$ and $\exp(0.05)>0.95>t_{2}>0$. Finally apply Lemma 8.3 under these bounds on $t_{1}$ and $t_{2}$. ∎
Remark 8.2.
We can extend our lower bounds of Lemma 8.3 and Cor. 8.1 on $\sigma_{r}(QH)$ to the case of any matrix $H$ of full rank $l$ if we decrease these bounds by a factor of $\sigma_{l}(H)$.
8.2 Output errors of Alg. 6.1 for a matrix with random singular space
Assumption 8.1.
Let $r<n\leq m$ (we can readily extend our study to the case where $m<n$). Fix two constant matrices
$\Sigma_{r}=\operatorname{diag}(\sigma_{j})_{j=1}^{r}~{\rm and}~\Sigma_{\perp}=\begin{bmatrix}\operatorname{diag}(\sigma_{j})_{j=r+1}^{n}\\ O_{m-n,n-r}\end{bmatrix}\in\mathbb{R}^{(m-r)\times(n-r)}$
such that $\sigma_{1}\geq\sigma_{2}\geq\dots\geq\sigma_{n}\geq 0$, and $\sigma_{r}>0$. Let $G$ be an $r\times n$ Gaussian matrix and let $Q\in\mathbb{R}^{n\times r}$ and $Q_{\perp}\in\mathbb{R}^{(n-r)\times r}$ be two matrices whose column sets make up orthonormal bases of the row space of $G$ and its orthogonal complement, respectively. Furthermore, let
$M=U\cdot\begin{bmatrix}\Sigma_{r}&0\\ 0&\Sigma_{\perp}\end{bmatrix}\cdot\begin{bmatrix}Q^{*}\\ Q_{\perp}^{*}\end{bmatrix}$ (8.1)
be SVD where the right singular space of $M$ is random and $U$ is a matrix with orthonormal columns.
Remark 8.3.
The matrix $G$ does not uniquely define the matrices $Q$, $Q_{\perp}$, and $U$ under Assumption 8.1 and hence does not uniquely define the matrix $M$, but this is immaterial for our analysis.
Theorem 8.2.
[Errors of Alg. 6.1 for an input with a random singular space.] Suppose that $G$ is an $r\times n$ Gaussian matrix, $H\in\mathbb{R}^{n\times l}$ is a constant matrix, $n>36r$, l>22(r1)l>22(r-1), $r\leq l<\min(m,n)$, and Alg. 6.1 applied to the matrix $M$ of (8.1) outputs two matrices $X$ and $Y$.
(i) If $H$ has orthonormal columns, then
[…]
with a probability no less than $1-\exp(-\frac{n}{72})-\exp(-\frac{l-r}{20})$.
(ii) If $H$ has full rank $l$, then
[…]
with a probability no less than $1-\exp(-\frac{n}{72})-\exp(-\frac{l-r}{20})$.
Proof.
We can assume that the matrices $G$ and $GH$ have full rank $r$ (see Thm. 8.1 and Remark 8.1). Consider SVD
$M=U\cdot\begin{bmatrix}\Sigma_{r}&0\\ 0&\Sigma_{\perp}\end{bmatrix}\cdot\begin{bmatrix}Q^{*}\\ Q_{\perp}^{*}\end{bmatrix}$,
write $C_{1}:=Q^{*}H$, apply Cor. 7.1, and deduce that
[…]
(i) Recall from Cor. 8.1 that
$\operatorname{Probability}\big{\{}\sigma_{r}(C_{1})=\sigma_{r}(Q^{*}H)\geq\frac{1}{4}\sqrt{l/n}\big{\}}>1-\exp\Big{(}-\frac{n}{72}\Big{)}-\exp\Big{(}-\frac{l-r}{20}\Big{)}.$ (8.2)
(ii) Let $H=U_{H}\Sigma_{H}V_{H}^{*}$ be a compact SVD. Then $\sigma_{r}(C_{1})\geq\sigma_{r}(Q^{*}U_{H})\sigma_{r}(H)$.
Similarly to (8.2) obtain that
$\operatorname{Probability}\big{\{}\sigma_{r}(C_{1})\geq\frac{1}{4}\sqrt{l/n}\cdot\sigma_{r}(H)\big{\}}$ $\geq\operatorname{Probability}\big{\{}\sigma_{r}(Q^{*}U_{H})\geq\frac{1}{4}\sqrt{l/n}\big{\}}$
$>1-\exp\Big{(}-\frac{n}{72}\Big{)}-\exp\Big{(}-\frac{l-r}{20}\Big{)}.$
Combine Thm. 7.1, equation $Q_{\perp}||_{2}=1$, and the bound $\sigma_{r}(C_{1})\geq\frac{1}{4}\sqrt{l/n}\cdot\sigma_{r}(H)$ and obtain
[…]
9 Numerical tests
In this section we cover our tests of dual superfast variants of Alg. 6.1. As in the tests of Part I we applied the standard normal distribution function randn of MATLAB to generate Gaussian matrices. The MATLAB function ”svd()” has been applied to calculate the $\epsilon$-rank for $\epsilon=10^{-6}$. The tests for Tables 9.19.3 have been performed on a 64-bit Windows machine with an Intel i5 dual-core 1.70 GHz processor by using custom programmed software in $C^{++}$ and compiled with LAPACK version 3.6.0 libraries.
9.1 Input matrices
We generated the following classes of input matrices $M$ for testing LRA algorithms.
Class I: $M=U_{M}\Sigma_{M}V_{M}^{*}$, where $U_{M}$ and $V_{M}$ are the Q factors of the thin QR orthogonalization of $n\times n$ Gaussian matrices, $\Sigma_{M}=\operatorname{diag}(\sigma_{j})_{j=1}^{n}$; $\sigma_{j}=1/j,~j=1,\dots,r$, $\sigma_{j}=10^{-10},~j=r+1,\dots,n$ (cf. [H02, Sec. 28.3]), and $n=256,512,1024$. (Hence $M||_{2}=1$ and $M^{+}||_{2}=10^{10}$.)
Class II: (i) The matrices $M$ of the discretized single-layer Laplacian operator of [39, Sec. 7.1]: $S\sigma](x)=c\int_{\Gamma_{1}}\log{|x-y|}\sigma(y)dy,x\in\Gamma_{2}$, for two circles $\Gamma_{1}=C(0,1)$ and $\Gamma_{2}=C(0,2)$ on the complex plane. We arrived at the matrices $M=(m_{ij})_{i,j=1}^{n}$, $m_{i,j}=c\int_{\Gamma_{1,j}}\log|2\omega^{i}-y|dy$ for a constant $c$, M=1||M||=1 and the arc $\Gamma_{1,j}$ of $\Gamma_{1}$ defined by the angles in the range $\frac{2j\pi}{n},\frac{2(j+1)\pi}{n}]$.
(ii) The matrices that approximate the inverse of a large sparse matrix obtained from a finite-difference operator of [39, Sec. 7.2].
Class III: The dense matrices of five families from Class II of Sec. 5.2.
We used $1024\times 1024$ SVD-generated input matrices of class I having numerical rank $r=32$, $400\times 400$ Laplacian input matrices of class II(i) having numerical rank $r=36$, $408\times 800$ matrices having numerical rank $r=145$ and representing finite-difference inputs of class II(ii), and $1000\times 1000$ matrices of class III, having numerical rank 4, 6, 10, 12, and 25 for the matrices of the five families wing, baart, foxgood, shaw, and gravity, respectively.
9.2 Five families of Ultrasparse sketch matrices $H$
We generated our $n\times(r+p)$ sketch matrices $H$ for random $p=1,2,\dots,21$ by using 3-ASPH, 3-APH (see the end of Appendix B.2), and Random permutation matrices. When the overestimation parameter $p$ was considerable, we actually computed LRA of numerical rank larger than $r$, and so LRA was frequently closer to an input matrix than the optimal rank-$r$ approximation. Accordingly, the output error norms in our tests ranged from about $10^{-4}$ to $10^{4}$ relative to the optimal errors.
We obtained every 3-APH and every 3-ASPH matrix by applying three Hadamard’s recursive steps (B.1) followed by random column permutation defined by random permutation of the integers from 1 to $n$ inclusive. While generating a 3-ASPH matrix we also applied random scaling with a diagonal matrix $D=\operatorname{diag}(d_{i})_{i=1}^{n}$ where we have chosen the values of independent identically distributed (iid) random variables $d_{i}$ sampled under the uniform probability distribution from the set $\{-4,-3,-2,-1,0,1,2,3,4\}$.
We used the following families of sketch matrices $H$: (0) Gaussian (for control), (1) sum of a 3-ASPH and a permutation matrix, (2) sum of a 3-ASPH and two permutation matrices, (3) sum of a 3-ASPH and three permutation matrices, (4) sum of a 3-APH and three permutation matrices, and (5) sum of a 3-APH and two permutation matrices.
9.3 Test results
Tables 9.19.6 display the average relative error norm $\frac{\|M-\tilde{M}\|_{2}}{\|M-M_{nrank}\|_{2}}$ in our tests repeated 100 times for each class of input matrices and each size of an input matrix and sketch matrix $H$ for Alg. 6.1 or for each size of an input matrix and a pair of left-hand and right-hand sketch matrices $F$ and $H$ for Alg. 6.3.
In all our tests we applied the sketch matrices of the six families of the previous subsection.
Tables 9.19.3 display the average relative error norm for the output of Alg. 6.1; in our tests it ranged from about $10^{-3}$ to $10^{1}$. The numbers in parentheses in the first line of Tables 9.2 and 9.3 show the numerical rank of input matrices.
Tables 9.49.6 display the average relative error norm for the output of Alg. 6.3 applied to the same input matrices from classes I–III as in our experiments for Alg. 6.1.
In these tests we used $n\times\ell$ and $\ell\times m$ sketch matrices for $\ell=r+p$ and $k=c\ell$ for $c=1,2,3$ and random $p=1,2,\dots,21$.
SVD-generated Matrices Laplacian Matrices Finite Difference Matrices
Family No. Mean Std Mean Std Mean Std
Family 0 4.52e+01 5.94e+01 6.81e-01 1.23e+$0$ 2.23e+$0$ 2.87e+$0$
Family 1 3.72e+01 4.59e+01 1.33e+$0$ 2.04e+$0$ 8.22e+$0$ 1.10e+01
Family 2 5.33e+01 6.83e+01 1.02e+$0$ 2.02e+$0$ 4.92e+$0$ 4.76e+$0$
Family 3 4.82e+01 4.36e+01 7.56e-01 1.47e+$0$ 4.82e+$0$ 5.73e+$0$
Family 4 4.68e+01 6.65e+01 7.85e-01 1.31e+$0$ 3.53e+$0$ 3.68e+$0$
Family 5 5.45e+01 6.23e+01 1.03e+$0$ 1.78e+$0$ 2.58e+$0$ 3.73e+$0$
Table 9.1: Relative error norms in tests for matrices of classes I and II
wing (4) baart (6)
Family No. Mean Std Mean Std
Family 0 1.07e-03 6.58e-03 2.17e-02 1.61e-01
Family 1 3.54e-03 1.39e-02 1.37e-02 6.97e-02
Family 2 4.74e-03 2.66e-02 1.99e-02 8.47e-02
Family 3 1.07e-03 5.69e-03 1.85e-02 8.74e-02
Family 4 4.29e-03 1.78e-02 8.58e-03 5.61e-02
Family 5 1.71e-03 1.23e-02 3.66e-03 2.38e-02
Table 9.2: Relative error norms for input matrices of class III (of San Jose University database)
foxgood (10) shaw (12) gravity (25)
Family No. Mean Std Mean Std Mean Std
Family 0 1.78e-01 4.43e-01 4.07e-02 1.84e-01 5.26e-01 1.24e+$0$
Family 1 1.63e+$0$ 3.43e+$0$ 8.68e-02 3.95e-01 3.00e-01 7.64e-01
Family 2 1.97e+$0$ 4.15e+$0$ 7.91e-02 4.24e-01 1.90e-01 5.25e-01
Family 3 1.10e+$0$ 2.25e+$0$ 4.50e-02 2.21e-01 3.63e-01 1.15e+$0$
Family 4 1.23e+$0$ 2.11e+$0$ 1.21e-01 5.44e-01 2.36e-01 5.65e-01
Family 5 1.08e+$0$ 2.32e+$0$ 1.31e-01 5.42e-01 2.66e-01 8.22e-01
Table 9.3: Relative error norms for input matrices of class III (of San Jose University database)
SVD-generated Matrices Laplacian Matrices Finite Difference Matrices
$k$ Class No. Mean Std Mean Std Mean Std
$\ell$ Family 0 2.43e+03 1.19e+04 1.28e+01 2.75e+01 9.67e+01 1.48e+02
Family 1 1.45e+04 9.00e+04 8.52e+03 8.48e+04 7.26e+03 2.47e+04
Family 2 4.66e+03 2.33e+04 3.08e+01 4.07e+01 3.80e+02 1.16e+03
Family 3 2.82e+03 9.47e+03 2.42e+01 3.21e+01 1.90e+02 3.90e+02
Family 4 3.15e+03 7.34e+03 2.71e+01 4.69e+01 1.83e+02 2.92e+02
Family 5 2.40e+03 6.76e+03 2.01e+01 3.56e+01 2.31e+02 5.33e+02
$2\ell$ Family 0 5.87e+01 5.59e+01 7.51e-01 1.33e+$0$ 3.17e+$0$ 3.89e+$0$
Family 1 7.91e+01 9.86e+01 3.57e+$0$ 7.07e+$0$ 1.55e+01 2.39e+01
Family 2 5.63e+01 3.93e+01 3.14e+$0$ 4.50e+$0$ 5.25e+$0$ 5.93e+$0$
Family 3 7.58e+01 8.58e+01 2.84e+$0$ 3.95e+$0$ 4.91e+$0$ 6.03e+$0$
Family 4 6.24e+01 4.54e+01 1.99e+$0$ 2.93e+$0$ 3.64e+$0$ 4.49e+$0$
Family 5 6.41e+01 6.12e+01 2.65e+$0$ 3.13e+$0$ 3.72e+$0$ 4.54e+$0$
$3\ell$ Family 0 9.29e+01 3.29e+02 8.33e-01 1.54e+$0$
Family 1 5.58e+01 4.20e+01 3.09e+$0$ 4.08e+$0$
Family 2 5.11e+01 4.94e+01 1.70e+$0$ 2.08e+$0$
Family 3 6.70e+01 8.27e+01 2.35e+$0$ 2.96e+$0$
Family 4 5.36e+01 5.74e+01 2.14e+$0$ 3.76e+$0$
Family 5 4.79e+01 4.58e+01 1.81e+$0$ 2.94e+$0$
Table 9.4: Relative error norms in tests for matrices of classes I and II
wing (4) baart (6)
$k$ Class No. Mean Std Mean Std
$\ell$ Family 0 1.70e-03 9.77e-03 4.55e+$0$ 4.47e+01
Family 1 3.58e+02 3.58e+03 1.42e-01 9.20e-01
Family 2 2.16e-01 2.10e+$0$ 1.10e-02 6.03e-02
Family 3 7.98e-04 7.22e-03 4.14e-03 3.41e-02
Family 4 5.29e-03 3.57e-02 2.22e+01 2.21e+02
Family 5 6.11e-02 5.65e-01 3.33e-02 1.30e-01
$2\ell$ Family 0 7.49e-04 5.09e-03 5.34e-02 2.19e-01
Family 1 4.74e-03 2.32e-02 2.14e-02 1.33e-01
Family 2 3.01e-02 2.34e-01 1.26e-01 7.86e-01
Family 3 2.25e-03 1.38e-02 5.91e-03 2.63e-02
Family 4 3.94e-03 2.54e-02 1.57e-02 6.71e-02
Family 5 2.95e-03 1.47e-02 1.58e-02 1.20e-01
$3\ell$ Family 0 4.59e-03 2.35e-02 1.50e-02 7.09e-02
Family 1 5.96e-03 2.82e-02 7.57e-03 4.84e-02
Family 2 1.74e-02 1.06e-01 6.69e-03 2.97e-02
Family 3 3.07e-03 3.07e-02 1.16e-02 5.16e-02
Family 4 2.57e-03 1.47e-02 2.35e-02 9.70e-02
Family 5 4.32e-03 2.70e-02 1.36e-02 5.73e-02
Table 9.5: Relative error norms for input matrices of class III (of San Jose University database)
foxgood (10) shaw (12) gravity (25)
$k$ Class No. Mean Std Mean Std Mean Std
$\ell$ Family 0 5.46e+$0$ 1.95e+01 8.20e-01 4.83e+$0$ 8.56e+$0$ 3.33e+01
Family 1 8.51e+03 1.88e+04 1.12e+$0$ 5.75e+$0$ 1.97e+01 1.00e+02
Family 2 5.35e+03 1.58e+04 1.93e-01 1.51e+$0$ 8.79e+$0$ 4.96e+01
Family 3 6.14e+03 1.74e+04 4.00e-01 1.90e+$0$ 7.07e+$0$ 2.45e+01
Family 4 1.15e+04 2.33e+04 2.95e-01 1.71e+$0$ 4.31e+01 3.80e+02
Family 5 7.11e+03 1.87e+04 2.18e-01 9.61e-01 6.34e+$0$ 2.59e+01
$2\ell$ Family 0 2.70e-01 7.03e-01 5.54e-02 2.62e-01 5.34e-01 1.59e+$0$
Family 1 5.24e+02 5.19e+03 4.67e-02 2.35e-01 1.38e+01 1.31e+02
Family 2 2.45e+$0$ 3.47e+$0$ 8.31e-02 6.37e-01 5.47e-01 1.69e+$0$
Family 3 2.43e+$0$ 3.74e+$0$ 1.24e-01 8.52e-01 5.10e-01 1.24e+$0$
Family 4 2.17e+$0$ 2.92e+$0$ 1.76e-01 8.76e-01 2.60e-01 7.38e-01
Family 5 2.10e+$0$ 3.34e+$0$ 1.26e-01 5.99e-01 5.68e-01 1.46e+$0$
$3\ell$ Family 0 2.62e-01 8.16e-01 4.49e-02 1.64e-01 4.59e-01 1.38e+$0$
Family 1 2.72e+$0$ 4.60e+$0$ 6.84e-02 3.43e-01 3.44e-01 8.60e-01
Family 2 2.42e+$0$ 3.92e+$0$ 8.26e-02 5.38e-01 6.89e-01 2.15e+$0$
Family 3 3.22e+02 3.20e+03 6.06e-02 2.95e-01 5.26e-01 1.17e+$0$
Family 4 1.91e+$0$ 3.36e+$0$ 6.61e-02 3.36e-01 6.19e-01 1.54e+$0$
Family 5 2.73e+$0$ 6.90e+$0$ 5.72e-02 2.39e-01 7.22e-01 1.59e+$0$
Table 9.6: Relative error norms for input matrices of class III (of San Jose University database)
10 Conclusions
[53] proposed and tested iterative refinement of LRA, and [28] has greatly advanced that study and incorporated the superfast algorithms of our present paper.
[27] proposed, analyzed, and tested superfast algorithms for matrix norm estimation.
[1] computed least squares approximation of multivariate functions (LSAMF) by means of reduction to the solution of Linear Least Squares Problem (LLSP) and further to C-A iterations for LRA but actually to any LRA algorithm, including our superfast ones.
[…]
Appendix
Appendix A Small families of hard inputs for superfast LRA
Any algorithm not involving all entries of an input matrix fails on the following input families.
Example A.1.
Let $\Delta_{i,j}$ denote an $m\times n$ matrix of rank 1 filled with 0s except for its $(i,j)$th entry filled with 1. The $mn$ such matrices $\{\Delta_{i,j}\}_{i,j=1}^{m,n}$ form a family of $\delta$-matrices. We also include the $m\times n$ null matrix $O_{m,n}$ filled with 0s into this family. Now fix any superfast algorithm; it does not access the $(i,j)$th entry of its input matrices for some pair of $i$ and $j$. Therefore, it outputs the same approximation of the matrices $\Delta_{i,j}$ and $O_{m,n}$, with an undetected error at least 1/2. Arrive at the same conclusion by applying the same argument to the set of $mn+1$ small-norm perturbations of the matrices of the above family and to the $mn+1$ sums of the latter matrices with any fixed $m\times n$ matrix of low rank. Finally, the same argument shows that a posteriori estimation of the output errors of an LRA algorithm applied to these input families cannot run superfast.
The example covers randomized LRA algorithms as well. Indeed, let an LRA algorithm miss an entry of an input matrix with a probability $p>0$. Apply this algorithm to two matrices of low rank whose difference at this entry is equal to a large constant $C$. Then with a probability $p$ the algorithm has error at least $C/2$ at this entry for a least one of these two matrices.
Appendix B LRA with dense and Abridged SRHT random sketch matrices
B.1 The generalized Nyström LRA algorithm (briefly)
Choose the generalized Nyström randomized algorithm of [48] to represent random sketching LRA algorithms. It modifies the algorithms of [18, Thm. 4.3, display 1] and [70] to improve numerical stability of rank-$r$ approximation of a matrix $M\in\mathbb{C}^{m\times n}$ for $r\leq n\leq m$. It first generates two random matrices $F\in\mathbb{C}^{k\times m}$ and $H\in\mathbb{C}^{n\times\ell}$, for $r\leq k\leq\ell\leq n$, then computes the sketches $FM$ and $MH$, and finally computes rank-$r$ approximation $\{X,Y,Z\}$ of $M$.
In the best studied case $F$ and $H$ are Gaussian random matrices, filled with independent standard Gaussian (normal) random variables; then for $k=2r+1$ and $\ell=2k$ the expected Frobenius error norm of the error matrix MXYZM-XYZ is within a factor of 2 from optimal (not exceeding the optimal one for rank-$\frac{k}{2}$ approximation). The algorithm involves $O((k+\ell)mn)$ flops, $mn$ entries of $M$, and about $km+n\ell+k\ell$ other scalars, overwriting some entries of $M$.
With SRHT or SRFT (rather than Gaussian) multipliers $F$ and $H$, one obtains LRA (with a little higher error probability) by using $O(mn\log(k\ell)))$ flops, $mn$ entries of $M$, and $mk\log(k)+n\ell\log(\ell)$ other scalars, also overwriting some entries of $M$.
Like the other random sketching LRA algorithms, this one is superfast except for the stage of the computation of the sketches $FM$ and $MH$ and can be readily modified to bound the relative error of the output LRA by any $\epsilon>0$, with estimated sketches’ size increasing fast as $\epsilon\mapsto 0$.
Remark B.1.
[Superfast LRA under the spectral norm.] The known algorithm supporting the same LRA error bounds under the spectral rather than Frobenius norm tend to be more involved (see [39, 25, 69]). E.g., the Power algorithm of [39] computes matrices $F(MM^{*})^{j}M$ and $(MM^{*})^{j}MH$, for random matrices $F$ and $H$, $j=0,1,\dots,t$, and $t$ of order of $\log(m+n)$. One can accelerate these computations by using sparse and/or structured matrices $F$ and $H$ but only for $j=0$ because multiplication by $MM^{*}$ destroys sparsity and structure.
B.2 Abridged SRHT matrices
With sparse subspace embedding [9, 12, 14], [71, Sec. 3.3], [46, Sec. 9] one obtains significant acceleration but still does not yield superfast algorithms. According to [40], such acceleration tends to make the accuracy of the output LRAs somewhat less reliable, although [14] partly overcomes this problem for incoherent matrices. One can multiply a matrix by Subsampled Randomized Hadamard or Fourier Transform (SRHT or SRFT) dense matrices towards incoherence [14], but this step is not superfast. By applying recipe of the end of Sec. 1.4, we devise Abridged SRHT multipliers, recall Thm. 8.2 for semi-empirical support for the output accuracy of the resulting LRA, and recall [57] for earlier formal and empirical study of their application to LRA.
We proceed by means of abridging the classical recursive processes of the generation of $n\times n$ SRHT matrices, obtained from the $n\times n$ dense matrices $H_{n}$ of Walsh-Hadamard transform for $n=2^{t}$ (cf. [44, Sec. 3.1]). Namely, we write
$H_{d,0}:=I_{n/2^{d}},~H_{d,i+1}:=\begin{pmatrix}H_{d,i}&H_{d,i}\\ H_{d,i}&-H_{d,i}\end{pmatrix}~{\rm for}~i=0,1,\dots,d-1.$ (B.1)
The $n\times n$ matrix $H_{d,d}$ is orthonormal up to scaling and has $2^{d}$ nonzero entries in every row and column. Now define $l\times n$ matrix $R$ of uniform random sampling of $l$ out of $n$ columns, the $n\times n$ diagonal matrix $D$ whose $n$ diagonal entries are independent random signs, i.e., random variables uniformly distributed on the pair $\{\pm 1\}$, and the $d$-Abridged subsampled randomized Hadamard transform (SRHT) matrix $\sqrt{2^{d}/l~}RH_{d,d}D$, with $2^{d}$ nonzero entries $\pm\sqrt{2^{d}/l}$ per column. We can multiply $M\in\mathbb{C}^{m\times n}$ by that matrix by involving $m2^{d}$ entries of $M$ and $(2^{d}-1)m$ flops.
Likewise, by sampling $k$ random rows of $H_{d,d}$ and multiplying the resulting matrix by $\pm\sqrt{2^{d}/k}D$ obtain $d$-Abridged $k\times m$ SRHT matrix having $2^{d}$ nonzero entries $\pm\sqrt{2^{d}/k}$ per row. We can multiply it by $M$ by involving $n2^{d}$ entries of $M$ and $(2^{d}-1)n$ flops, where a flop amounts to addition, subtraction, or multiplications by $\sqrt{2^{d}/l}$ or $\sqrt{2^{d}/k}$. In particular this means $8n$ (resp. $8m$) entries of $M$ and $7n$ (resp. $7m$) flops in our numerical tests, where we always fix $d=3$.
A $t$-Abridged SRHT matrix turns into an SRHT matrix.
Table B.1 displays the computational cost of randomized rank-$r$ approximation by means of the algorithm of [48], when it uses Gaussian, SRHT and $d$-Abridged SRHT multipliers $F$ and $H$.
Multipliers $F$ and $H$ Input entries involved Other entries involved Flops involved
Gaussian mnmn $km+n\ell+kl$ $O((k+l)mn)$
SRHT mnmn $O(mk\log(k)+nl\log(l))$ $O(mn\log(kl))$
$d$-Abridged SRHT $2^{d}(m+n)$ $2^{d}(k+l)$ $(2^{d}-1)(m+n)$
Table B.1: Computational cost of the generalized Nyström algorithm of [48] with Gaussian, SRHT, and $d$-Abridged SRHT multipliers $F$ and $H$.
We also refer to $d$-abridged SRHT matrices as $d$-Abridged Scaled and Permuted Hadamard $d$-ASPH matrices and call them $d$-APH matrices if random scaling is omitted, that is, if the multiplier $D$, filled with $\pm 1$, is replaced with the identity matrix.
Appendix C The spectral norms of a Gaussian matrix and its pseudo inverse
Hereafter $\Gamma(x)=\int_{0}^{\infty}\exp(-t)t^{x-1}dt$ denotes the Gamma function; $\nu_{p,q}$ and $\nu^{+}_{p,q}$ denote the random variables representing the spectral norms of a $p\times q$ Gaussian random matrix and its Moore-Penrose pseudo inverse, respectively.
Theorem C.1.
[Spectral norms of a Gaussian matrix. See [21, Thm. II.7].]
[…]
Theorem C.2.
[Spectral norms of the pseudo inverse of a Gaussian matrix.]
(i) Probability $\{\nu_{m,n}^{+}\geq m/x^{2}\}<\frac{x^{m-n+1}}{\Gamma(m-n+2)}$ for $m\geq n\geq 2$ and all positive $x$,
(ii) Probability $\{\nu_{m,n}^{+}\geq t\frac{e\sqrt{m}}{m-n+1}\}\leq t^{n-m}$ for all $t\geq 1$ provided that $m\geq 4$,
(iii) $\mathbb{E}(\nu^{+}_{m,n})\leq\frac{e\sqrt{m}}{m-n}$ provided that $m\geq n+2\geq 4$,
Proof.
See [10, Proof of Lemma 4.1] for claim (i), [39, Prop. 10.4 and Eqns. (10.3) and (10.4)] for claims (ii) and (iii), and [65, Thm. 3.3] for claim (iv). ∎
Thm. C.2 implies reasonable probabilistic upper bounds on the norm $\nu_{m,n}^{+}$, even where the integer |mn||m-n| is close to 0; whp the upper bounds of Thm. C.2 on the norm $\nu^{+}_{m,n}$ decrease very fast as the difference |mn||m-n| grows from 1.
Acknowledgements: Our work has been supported by NSF Grants CCF–1116736, CCF–1563942 and CCF–1733834 and PSC CUNY Awards 66720-00 54 and 68452-00 56. We are also grateful to E. E. Tyrtyshnikov for the challenge of formally supporting empirical power of C–A iterations, to N. L. Zamarashkin for his comments on [55], and to S. A. Goreinov, I. V. Oseledets, A. Osinsky, E. E. Tyrtyshnikov, and N. L. Zamarashkin for reprints and pointers to relevant bibliography.