### v3

# Novel Insights into Low Rank Approximation at Sublinear Cost

Abstract
Low rank approximation of a matrix (LRA) is a highly important area of Numerical Linear and Multilinear Algebra and Data Mining and Analysis. One can operate with an LRA superfast – by using much fewer memory cells and flops than an input matrix has entries.111“Flop” stands for “floating point operation”. Can we, however, compute an LRA of a matrix superfast? YES and NO. For worst case inputs, any LRA algorithm fails miserably unless it involves all input entries, but in computational practice worst case inputs seem to appear rarely, and accurate LRA are routinely computed superfast for large and important classes of matrices, in particular in the memory efficient form of CUR, widely used in data analysis. We advance formal study of this YES and NO coexistence by proving novel universal upper bounds on the spectral output error norms of all CUR LRA algorithms and, under a fixed probabilistic structure in the space of input matrices, on both spectral and Frobenius error norms of nearly all sketching LRA algorithms. These bounds imply that superfast LRA algorithms of the two kinds fail miserably only for a very narrow input class. Furthermore, in our numerical tests such superfast algorithms were consistently much more accurate than our upper estimates ensure and usually were reasonably close to optimal.
Key Words:
Low-Rank Approximation, CUR LRA, Sublinear algorithms.
2020 Math. Subject Classification:
65Y20, 65F55, 68Q25, 68W20
1 Introduction
1.1. Superfast LRA CUR LRA: background. LRA of a matrix is among the most fundamental problems of Numerical Linear and Multilinear Algebra and Data Mining and Analysis, with applications ranging from machine learning theory and neural networks to term document data and DNA SNP data (see surveys [30, 37, 59]). For example, matrices that represent Big Data (e.g., unfolding matrices of multidimensional tensors) tend to be so immense that only a tiny fraction of their entries fits primary memory of a computer, but quite typically they admit LRA, that is, can be closely approximated with the products of a few matrices of smaller sizes [62]. Such products can be multiplied by a vector superfast, aka at sublinear cost, – by using much fewer memory cells and flops than the matrix has entries. [^2] Superfast algorithms have been studied long and extensively for structured matrices having small displacement rank (see more in Sec. 1.4.1), and we extend this concept to computations of LRA. This implies dramatic acceleration of various popular matrix computations reduced to that basic operation, but can we compute an LRA of a matrix superfast or at least quasi superfast, that is, by involving only a small fraction of all entries of an input matrix? YES and NO.
NO, because on the worst case inputs and even on the small matrix families of our Example A.1 any algorithm fails miserably for computing accurate LRA and even for a posteriori estimation of its accuracy unless it involves all entries of an input matrix.
YES, because in computational practice worst case inputs seem to appear rarely, and accurate LRAs are routinely computed superfast for a large and important class of matrices, in particular by means of the popular Cross-Approximation (C–A) iterations, specializing the Alternating Directions Implicit (ADI) method [54] to LRA and recalled in Sec. 3.2.
Their output LRAs are memory efficient products CUR where the factors $C$ and $R$ are made up of $l$ columns and $k$ rows of an input matrix, respectively, the factor $U$ has size $k\times l$,
$r\leq k\leq m,~r\leq l\leq n,~kl\ll mn,$ (1.1)
and $r$ denotes a target rank,333“In practice, the target rank is rarely known in advance. $\dots$ LRA algorithms are usually implemented in an adaptive fashion” – with submatrices $R$ and $C$ recursively expanded “until the error norm satisfies the desired tolerance” [60, Sec. 2.4]. presumed to be low. The $k\times l$ submatix $G$ shared by the factors $C$ and $R$ is said to be CUR generator. Given $M$ and $G$, the factors $C$ and $R$ are immediately available, and one obtains canonical CUR by computing the factor $U$ in about $kl$ scalars and $O(kl\min\{k,l\})$ flops (see (2.3)). This computation is quasi superfast for $kl\ll mn$ and superfast for $kl\min\{k,l\}\ll mn$.
In [67] Zamarashkin and Osinsky proved that, for any $m\times n$ matrix $M$ and any positive integer $k\leq\min\{m,n\}$, there exists a CUR LRA defined by a $k\times k$ generator and optimal up to a factor of $k+1$ under the Frobenius matrix norm. In [13] Cortinovis and Kressner computed such a CUR at deterministic cost $O(kmn^{3})$ for $m\geq n$ by means of linking the task to subset selection. Superfast computation of a good generator, however, is a challenge, and the error norms of CUR LRAs of $M$ are unbounded for ill-conditioned generators $G$.
1.2. Our contribution. We extract our Part I and Part II from preprints [47, 48], respectively (see further comments on this in Sec. 1.4.3), and point out applications of LRA to superfast matrix norm estimation and least squares approximation of multivariate functions in Secs. 3.3 and 11.
1.2.1. Part I: the study of CUR LRA. According to our Main Result 1 (Thm. 4.1), any CUR LRA (no matter how computed) has spectral output error norm $||\cdot||$ within a factor of $(v+1)^{2}$ from optimal for
[…]
This universal upper bound implies that any CUR LRA can fail miserably only on a narrow class of inputs. In our tests various known and novel superfast CUR LRA algorithms were consistently much more accurate than this upper bound ensures and usually have spectral output error norms within a factor of 30 from optima for $n\times n$ inputs and $n=256,512,1024$ (see Table 6.1).
Furthermore, instead of customary combinations of C-A iterations Strong Rank Revealing (SRR) QR or LU factorization of [19, 41], we tested their combination with random sampling algorithms of [16], and in our tests the output accuracy of this superfast CUR LRA was almost always nearly as accurate as in the case of the algorithms of [16] themselves, which run at superlinear cost and are expected to be near-optimal under the Forbenius norm (see Tables 6.2 and 6.3).
1.2.2. Part II: sketching LRA algorithms. We provide another insight into this subject by proving a little stronger error bound for a large class of superfast LRA algorithms. First recall that with a high probability (whp) random sketching LRA algorithms of [64, 60] compute near-optimal rank-$r$ approximation [^4] Here and hereafter “rank-$r$ approximation” means “approximation of rank at most rr”. of an $m\times n$ matrix $M$ under the Frobenius norm by using auxiliary matrices (sketches) $FM$ and/or $MH$ for proper $k\times m$ and/or $n\times l$ random multipliers $F$ and/or $H$, respectively, $\max\{k,l\}=r+p$, $k,l,m,n,r$ satisfying (1.1), and a positive oversampling (overestimation) parameter $p\ll\min\{m,n\}$. Usually these algorithms output LRA but not CUR LRA (see [16] for an important exception), but one can extend any LRA to CUR LRA – at the price of increasing the spectral norm bounds on the output errors by a factor of $\sqrt{1+(n-r)r)}$ (see [30, Secs. 3.3.2 and 3.3.3]).
Apart from the stage of the computation of sketches, random sketching algorithms of [30] (resp. of [60]) and their extensions in our Sec. 7 run superfast where $l^{2}\ll n$ (resp. $l^{2}\ll n$ and $k^{2}\ll m$) and quasi superfast where $l\ll n$ (resp. $l\ll n$ and $k\ll m$). We compute the sketches superfast as well by choosing Ultrasparse multipliers $F$ and/or $H$ (see Appendix B.2). This way formal support for near-optimality of LRA whp is lost, but next we specify our weaker but meaningful alternative.
To simplify our presentation, we restrict it to real matrices throughout, with $M^{*}$ denoting transpose rather than Hermitian transpose of $M$, and to the random sketching LRA that only uses sketches $MH$, as in [30], while in Sec. 8.2 we extend our results to algorithms that use both sketches $FM$ and $MH$.
One can explore a variety of ways towards our goal. Maybe the simplest way is to randomly sample $l$ columns of the matrix $M$ or equivalently to let $H$ be the $n\times l$ leftmost submatrix of a random $n\times n$ permutation matrix under a fixed randomization model; in this case only $nl$ elements of $M$ are involved, and the calculation itself is given ”for free”. We, however, consider the more general class of all sketches $MH$ that can be computed superfast. Example A.1 tells us that we cannot yield accurate LRA for worst case inputs $M$ superfast, but in our numerical experiments with Ultrasparse multipliers $F$ and $H$ we have consistently output quite accurate LRA for a large class of matrices $M$.
Towards explaining these test results and constructing a theory, we introduce a probabilistic structure in the space of input matrices $M$ rather than in the space of multipliers $H$, thus defining dual sketching in comparison to customary primal one, where matrix $M$ is fixed and multiplier $H$ is random.
[…]
In Sec. 9.3 we study dual LRAs under another choice of probabilistic structure, which may be more relevant to LRA computation in the real world: we fix a multiplier $H$ with orthonormal columns and let $M:=A\Sigma B+E$ for a perturbation matrix $E$ and for $A\Sigma B$ being a random pseudo SVD of a rank-$r$ matrix of size $m\times n$, said to be a two-sided factor-Gaussian (with expected rank $r$). Namely, we let $\Sigma$ be any $r\times r$ matrix of full rank $r$ (e.g., a diagonal matrix with $r$ positive diagonal entries, as in SVD) and let $A$ and $B$ be scaled Gaussian matrices, rather than matrices of singular vectors of SVD. [^5] By saying “LRA” we assume that $r$ is much smaller than $\min\{m,n\}$; then we can motivate the above definition of random pseudo SVD by recalling (e.g., from [18, Theorem 7.3] or [53]) that $\frac{1}{\sqrt{k}}G$ is close to a matrix with orthonormal columns whp for $r\ll k$ and an $n\times r$ Gaussian matrix $G$. Most of our study (in particular Theorem 9.1 and our Main Result 3, outlined below) applies to a more general class of perturbed right factor-Gaussian matrices, of the form $AG+E$ where $G$ is a $r\times n$ Gaussian matrix and $A$ is an $m\times r$ matrix of full rank $r$; transposition extends the study to perturbed left factor-Gaussian matrices.
For a matrix $M$ with a certain gap in the spectrum of its singular values [^6] This assumption is restrictive but does not exclude highly important kernel matrices [11, 65]. we prove that our superfast sketching algorithms output close rank-$r$ approximations whp on the defined probability space: namely, our bound on the output errors under this randomization model only increases from a factor of $\sqrt{1+16~n/l}$ versus the optimum to a factor of $\sqrt{1+100~n/l}$, and this is essentially Thm. 9.2 – our Main Result 3. Over the same class of factor-Gaussian matrices we also strengthen the bound of our Main Result 1 whp (see Remark 5.1).
In our tests we sampled random multipliers $F$ and $H$, even though our Main Results 2 and 3 (Thms. 9.1 and 9.2) also hold where these multipliers are fixed.
1.3. Quality of superfast dual LRA. Can we decrease the upper bounds of Thms. 9.1 and 9.2 if we choose multipliers $H$ according to a proper heuristic policy? Such a policy can be shaped empirically, based on testing various candidate multipliers $H$, but here is our semi-heuristic considerations. The best-studied LRA algorithms use Gaussian multipliers $F$ and $H$, output near-optimal LRA whp, but run at superlinear cost. With sketches obtained by using sparse subspace embedding [10, 12], [61, Sec. 3.3], [37, Sec. 9], one devises LRA algorithms running at linear computational cost. According to [32], such acceleration tends to make the output accuracy somewhat less reliable. This problem is partly overcome for incoherent matrices, whose all entries have comparable magnitude (see [12] for this study and formal definition of incoherence). Multiplication by the matrix of Subsampled Randomized Hadamard or Fourier Transform (SRHT or SRFT) tends to make a matrix $M$ incoherent [12] but at superlinear cost.
Now, for our heuristic recipe, we devise $d$-Abridged SRHT multipliers in Appendix B.2.777Abridged SRFT matrices can be defined similarly to Abridged SRHT matrices and empirically have similar power (see [46, Sec. 13.6 and Appendix B]); unlike real Abridged SRHT matrices they involve nonreal roots of unity. For $d=\log_{2}(n)$ they turn into $n\times l$ SRHT multipliers, but for small positive integers $d$ (we use $d=3$ in our tests) they are Ultrasparse and can be multiplied by a dense matrix superfast. Furthermore, they have orthonormal columns, and so Thm. 9.1 provides dual probabilistic motivation for this otherwise heuristic choice of multipliers. In Sec. 10 we provide its empirical support by testing our superfast extensions of the algorithms of [30, 60]. Empirical benefits of $d$-Abridged SRHT multipliers were observed in [48] and apparently are still a novelty. In extensive numerical tests in [21] for $d=3$ they were at least as efficient as Gaussian most of the time although caused blow up of the output errors of LRA for a narrow class of inputs.
With this reservation we can use $d$-abridged SRHT multipliers instead of Gaussian towards additional support of our dual model in Appendix D. There we prove that pre- and/or post- multiplication by Gaussian matrices turns any matrix admitting LRA into a small-norm perturbation of a factor-Gaussian matrix. This multiplication becomes superfast if we substitute $3$-abridged SRHT multipliers for Gaussian.
At the end of Sec. 1.5 we comment on goals and Abridged SRHT by-product of our numerical tests.
1.4. Related work. We recall two groups of the known studies of superfast LRA.
1.4.1. Formally supported superfast LRA. Superfast algorithms have been long studied for data sparse matrices – defined by small numbers of parameters such as ones with displacement structure (cf. [31, 5, 42, 66, 43] and the bibliography therein), but some of more recent superfast algorithms output LRAs of the families of data dense matrix families, defined by order of $mn$ parameters in the space of $m\times n$ matrices. In this subsection such families are specified formally, although in some cases in terms of parameters not available superfast.
The randomized superfast algorithms of [38, 11] and [7] compute LRA for $n\times n$ Symmetric Positive Semidefinite (SPSD) and distance matrices, respectively. The superfast algorithms of [38, 7] are near-optimal whp under the Frobenius norm, while the algorithm of [11], based on random pivoted Cholesky factorization, outputs oversampled LRA of size $k\times k$ for $k>r/\epsilon$ and hence is not superfast unless $\epsilon\gg n/r$; the algorithm, however, is highly competitive in practice, particularly for SPSD kernel matrices, unlike that of [38] (see [2, 11, 65] and the references therein for kernel matrices and their applications).
The sublinear cost bounds of [7, 38, 11] do not cover superlinear cost of a posteriori estimation of their output error norms and correctness verification. The superfast deterministic algorithm of [33, Part III], however, computes CUR LRA of an $n\times n$ SPSD matrix with both spectral and Frobenius error norms within a factor of $n$ from optimal and as by-product, at no additional cost, estimates the output error norm, verifying correctness.
Chiu and Demanet in [9] defined CUR generator by sampling rows and columns of an input matrix $M$ uniformly at random and proved that this algorithm, which we call Primitive, is quite accurate under the Frobenius norm whp provided that it is directed by a proper “regularization parameter” $\delta$ and that there exists an accurate LRA $XYZ\approx M$, e.g., defined by the SVD of $M$, such that the matrices $X$ and $Z^{*}$ have orthonormal columns and are incoherent.
The algorithm is superfast except for the computation of proper regularization parameter $\delta$, and a numerical example in [9, Sec. 4.1] shows that without this regularization, the approximation error can blow up. Likewise, in our tests in Sec. 6 the output error norm of that algorithm performed without regularization exceeded the norms of all other tested CUR LRA algorithms by factors ranging from 100 to 5,000 in the case of $n\times n$ inputs for $n=256,512,1024$.
In [9, Sec. 3.2] superfast Interpolative Decomposition of [30, Sec. 3.2.3] based on QR factorization with column pivoting is added to that algorithm, and then the authors proved that their estimates for the output accuracy still hold where just one of the factors $X$ or $Z^{*}$ has orthonormal columns and is incoherent.
Cortinovis and Ying in [15] make a little weaker assumptions about the matrices $X$ and $Z^{*}$ – they allow some of their column vectors to be sparse rather than incoherent provided that CUR LRA algorithm incorporates another variant of the Interpolative Decomposition, used earlier in [34, 65]; we outline it in the next subsection. CUR LRA algorithm of [15] was quite accurate in the tests in [15] but, as in [9], is not superfast – it relies on non-canonical transition from CUR generator to CUR LRA based on (2.2).
Although the algorithms of [9, 15] are not superfast, analysis of their output accuracy is still interesting, but the provision of incoherence of the factors of an unknown LRA of $M$ does not hold for a large class of inputs and generally cannot be verified superfast for a fixed $M$.
Our Main Results 1 and 2 (Thms. 4.1 and 9.1) ensure weaker output accuracy than [9, 15] but are universal – they apply to nearly all input matrices and either to all CUR LRAs or to all sketching LRA algorithms, respectively.
1.4.2. Heuristic superfast LRA. Next we recall some superfast LRA algorithms, whose output accuracy was empirically stronger than our Main Results ensure, similarly to what we observed in our tests.
C-A iterations is the most popular example of such algorithms. For two decades of their worldwide application they have been consistently computing accurate CUR LRA for a large class of matrices (see [23, 3, 40, 1], and the bibliography therein).
The heuristic superfast [34, Alg. 2.2] is much less known but important as well. Its output CUR LRA is quite accurate for a large class of unsymmetric $n\times n$ matrices $M$ according to the test results in [34].
For a target rank $r$ this algorithm seeks column and row sets that define the factors $C$ and $R$ of CUR. Initially the two sets are empty and then are recursively increased “a few times” based on nontrivial extension of the algorithm of [9, Sec. 3].
Namely, one first uniformly samples $q>r$ rows of $M$ to obtain a $q\times n$ submatrix, then applies to this submatrix QR factorization with pivoting as in [9], and finally appends the first $r$ column indexes defined by this factorization to the current set of column indexes.
One expands the current set of row indexes by applying the same algorithm to the transpose $M^{*}$.
When this recursive computation of two factors $C$ and $R$ of the CUR of $M$ is completed, a factor $U$ is computed by means of solving a generalized LLSP defined by $M$, $C$ and $R$. This heuristic algorithm involves order of $n^{3/2}$ memory cells and flops, and the authors specify a direction to decreasing that bound to order of $n\log(n)$ but admit: “We have not been able to quantify the error and success probability rigorously for this procedure at this point.”
Instead of QR factorization with column pivoting, Jianlin Xia in [65] uses SRR QR factorization of [19] and calls this technique “progressive alternating direction pivoting”. He combines it with his additional techniques, in particular, randomized error estimation. His algorithms, “numerically shown to have nearly linear complexity” in $m+n$, have consistently output accurate LRAs of real world $m\times n$ kernel matrices $M$ in his extensive numerical tests. He elaborates upon various techniques used for devising his algorithms, but not as much upon their formal support and estimates for their output accuracy and complexity. He admits: “Although a fully rigorous justification of the accuracy is lacking, we give different perspectives to motivate and support the ideas”, but it is still a challenge to specify formally the large subclass of kernel matrices for which his algorithms output accurate LRA. He cites [33, 49] as his close predecessors, but neither his algorithms nor [34, Alg. 2.2] significantly overlap with our current ones for CUR LRA and involve no random sketching at all.
1.4.3. Our old preprints. In Parts I and II we extend the unpublished preprints [45, 46, 47, 48], which cite [50, 51, 52] as their predecessors, where matrix computations under the dual probabilistic model have been formally studied, apparently for the first time. [^8] Such a model was implicit in classical initialization of the Power method with multiplication by a coordinate vector but has not been formally studied or even introduced. The cited preprints stated motivation for further study of superfast LRA as their major goal (which we share) and seemed to be too unorthodox when they appeared, but lately, encouraged by the success of [65, 15], we revisited preprints [47, 48] and extracted from them our Parts I and II, respectively. To present our work in proper historical context, we keep the results of the numerical tests of 2019 intact and do not extend them, e.g., to the numerically stable algorithm of [39], even though this could have strengthened their performance.
1.5. Organization of our paper. We recall some background material on LRA, Gaussain and factor-Gaussian matrices, and CUR LRA in the next two subsections and Sec. 2, respectively. In Part I (Secs. 26) we study computation of CUR LRA. In Sec. 3 we briefly comment on C-A iterations and specify some more rudimentary superfast CUR LRA. In Sec. 4 we estimate output errors of any CUR LRA, thus proving our Main Result 1. In Sec. 5 we a little strengthen it in the case of a perturbed factor-Gaussian input matrix. We devote Sec. 6 to our numerical experiments for the algorithms of Part I. In Part II (Secs. 710) we study random sketching algorithms for LRA. We recall and a little extend some known ones in Sec. 7 and recall their deterministic output error bounds in Sec. 8. In Sec. 9 we prove our Main Results 2 and 3, on dual output error bounds. In Sec. 10 we cover numerical tests of the algorithms of Part II. We devote Sec. 11 to conclusions. In Appendix A we specify some small families of matrices on which any LRA fails miserably unless it involves all entries of $M$. In Appendix B we first comment on some random sketching LRA and then define Abridged SRHT Ultrasparse matrices. In Appendix C we recall the estimates for the norms of a Gaussian matrix and its pseudo inverse. In Appendix D we prove that pre- and post-multiplication by Gaussian matrices turns any rank-$r$ matrix into a factor-Gaussian matrix having expected rank $r$. In Appendix E we prove necessary and sufficient condition for having CUR decomposition of a matrix.
In Secs. 3 and 7 we describe large families of LRA algorithms covered by our Main Results. In Secs. 6 and 10 we tested numerically some specified algorithms from these families, which supported our Main Results empirically. This was the only purpose of our numerical tests, but the observation of potential benefits of using $d$-Abridged SRHT multipliers has come as a by-product of that study.
1.6. Some background for LRA. In this subsection, we write $|||\cdot|||$ to unify the spectral norm $||\cdot||$ and the Frobenius norm $||\cdot||_{F}$, by following [30, Thm. 9.1].
Hereafter $\mathbb{R}^{p\times q}$ denotes the class of $p\times q$ real matrices.
For positive integers $k,l,m,n,r$ satisfying (1.1), the triplet of matrices
$X\in\mathbb{R}^{m\times k},~Y\in\mathbb{R}^{k\times\ell},~Z\in\mathbb{R}^{n\times\ell},~r\leq\min\{k,\ell\},$ (1.3)
as well as the product $XYZ^{*}$ represent a rank-$r$ approximation of a matrix $M\in\mathbb{C}^{m\times n}$ with the error matrix $E:=M-~XYZ^{*}$. One can replace the triplet $\{X,Y,Z^{*}\}$ with the pair $\{XY,Z^{*}\}$ of $m\times\ell$ and $\ell\times n$ matrices or $\{X,YZ^{*}\}$ of $m\times k$ and $k\times n$ matrices.
An important 3-factor LRA of $M$ is its $r$-top SVD $M_{r}=U_{r}\Sigma_{r}V_{r}^{*}$ for a diagonal matrix $\Sigma_{r}=\operatorname{diag}(\sigma_{j})_{j=1}^{r}$ of the $r$ largest singular values of $M$ and two matrices $U_{r}$ and $V_{r}$ of the $r$ associated left and right singular vectors, respectively.
$M_{r}$ is said to be the $r$-truncation of $M$. One can obtain it from a matrix $M$ by setting to 0 all its singular values but the $r$ top (largest) ones. $r$-top SVD of $M$ is compact SVD of $M_{r}$.
$M^{+}$ denotes the Moore–Penrose pseudo inverse of $M$.
Theorem 1.1.
(Eckart-Young-Mirsky Theorem, see [22, Thm. 2.4.8].) It holds that
[…]
under both spectral and Frobenius norms: $\tau_{r+1}(M)=\sigma_{r+1}(M)$ under the spectral norm, and $\tau_{r+1}(M)=\sigma_{F,r+1}(M):=\sqrt{\sum_{j>r}\sigma_{j}^{2}(M)}$ under the Frobenius norm.
Lemma 1.1.
(The norm of the pseudo inverse of a matrix product, see, e.g., [27].) Suppose that $A\in\mathbb{R}^{k\times r}$, $B\in\mathbb{R}^{r\times l}$ and the matrices $A$ and $B$ have full rank $r\leq\min\{k,l\}$. Then
[…]
Theorem 1.2.
[The impact of a perturbation of a matrix on its singular values (see [22, Cor. 8.6.2]).] For a pair of $m\times n$} matrices $M$ and $M+E$ it holds that
[…]
Lemma 1.2.
(The impact of a perturbation of a matrix on its singular space, adapted from [63], [55, Thm. 6.4], [26, Thm. 1].) Let $M$ be an $m\times n$ matrix of rank $r<\min(m,n)$ where
$M=\begin{bmatrix}U_{r}&U_{\perp}\end{bmatrix}~\begin{bmatrix}\Sigma_{r}&0\\ 0&0\end{bmatrix}~\begin{bmatrix}V_{r}^{T}\\ V_{\perp}^{T}\end{bmatrix}$
is its SVD, and let $E$ be a perturbation matrix such that
[…]
Remark 1.1.
Matrix $\tilde{V}$ above does not necessarily have orthogonal columns, but one can readily prove that the matrix $(V_{r}+V_{\perp}P)(I_{r}+P^{T}P)^{-1/2}$ has orthonormal columns, i.e., $(I_{r}+P^{T}P)^{-1/2}$ normalizes $\tilde{V}$ (see [63, 55, 26]).
1.7. Background on Gaussian and factor-Gaussian matrices. Constant matrices are filled with constants, unlike random matrices, filled with random variables.
Theorem 1.3.
[Non-degeneration of a Gaussian matrix.] Suppose that $M\in\mathbb{R}^{p\times q}$ is a constant matrix, $r\leq\operatorname{rank}(M)$, and $F$ and $H$ are $r\times p$ and $q\times r$ independent Gaussian matrices, respectively. Then the matrices $F$, $H$, FMFM, and $MH$ have full rank $r$ with probability 1.
Proof.
Rank deficiency of matrices $F$, $H$, FMFM, and $MH$ is equivalent to turning into 0 the determinants $\det(FF^{*})$, $\det(H^{*}H)$, $\det(FM(FM)^{*})$, and $\det((MH^{*})MH)$, respectively. The claim follows because these equations define algebraic varieties of lower dimension in the linear spaces of the entries, considered independent variables (cf., e.g., [6, Prop. 1]). ∎
Remark 1.2.
Events that occur with probability 0 are immaterial for our probability estimates, and hereafter we say that a matrix has rank $r$ if it is has rank $r$ with probability 1.
Lemma 1.3.
[Orthogonal Invariance.] [58, Theorem 3.2.1]. Suppose that $G$ is an $m\times n$ Gaussian matrix, $k\leq\min\{m,n\}$ is a positive integer, and $S\in\mathbb{R}^{k\times m}$ and $T\in\mathbb{R}^{n\times k}$ are constant matrices, having orthonormal rows and columns, respectively. Then $SG$ and $GT$ are random matrices having distribution of $k\times n$ and $m\times k$ Gaussian random matrices, respectively.
Definition 1.1.
[Factor-Gaussian matrices.] Let $A\in\mathbb{R}^{m\times r}$, $B\in\mathbb{R}^{r\times n}$, and $C\in\mathbb{R}^{r\times r}$ be three constant well-conditioned matrices of full rank $r<\min\{m,n\}$. Let $G_{1}$ and $G_{2}$ be $m\times r$ and $r\times n$ independent Gaussian matrices, respectively. Then $G_{1}B$, $AG_{2}$, and $G_{1}CG_{2}$ are left, right, and two-sided factor-Gaussian matrices of rank $r$, respectively.
Theorem 1.4.
The distribution of two-sided $m\times n$ factor-Gaussian matrices $G_{m,r}CG_{r,n}$ does not change if in its definition we replace the factor $C$ by an appropriate diagonal matrix $\Sigma=(\sigma_{j})_{j=1}^{r}$ such that $\sigma_{1}\geq\sigma_{2}\geq\dots\geq\sigma_{r}>0$.
Proof.
Let $C=U_{C}\Sigma_{C}V_{C}^{*}$ be SVD. Then $A=G_{m,r}U_{C}$ and $B=V_{C}^{*}G_{r,n}$ have distributions of independent $m\times r$ and $r\times n$ Gaussian matrices, respectively, by virtue of Lemma 1.3. Hence $G_{m,r}CG_{r,n}$ has the same distribution as $A\Sigma_{C}B$. ∎
PART I: COMPUTATION OF CUR LRA
2 CUR decomposition and CUR LRA
For two sets $\mathcal{I}\subseteq\{1,\dots,m\}$ and $\mathcal{J}\subseteq\{1,\dots,n\}$, define the submatrices
$M_{\mathcal{I},:}:=(m_{i,j})_{i\in\mathcal{I};j=1,\dots,n},M_{:,\mathcal{J}}:=(m_{i,j})_{i=1,\dots,m;j\in\mathcal{J}},~{\rm and}~M_{\mathcal{I},\mathcal{J}}:=(m_{i,j})_{i\in\mathcal{I};j\in\mathcal{J}}.$
Given an $m\times n$ matrix $M$ of rank $r$ and its nonsingular $r\times r$ submatrix $G=M_{\mathcal{I},\mathcal{J}}$, one can readily verify that $M=M^{\prime}$ for
$M^{\prime}=CUR,~C=M_{:,\mathcal{J}},~U=G^{-1},~G=M_{\mathcal{I},\mathcal{J}},~{\rm and}~R=M_{\mathcal{I},:}.$ (2.1)
We call $G$ the generator and $U$ the nucleus of CUR decomposition of $M$ where $X=C$, $Y=U$, $Z=R^{*}$, and $E=O$).
CUR decomposition is extended to CUR approximation of a matrix $M$, although the approximation $M^{\prime}\approx M$ for $M^{\prime}$ of (2.1) tends to be poor where the generator $G$ is ill-conditioned.
Remark 2.1.
The pioneering papers of 1995-2001 (cf. [27, 25] and the references in [40]) define CGR approximations having nuclei $G$; “G” can stand, say, for “germ”. We use the acronym CUR, which is more customary in the West. “U” can stand, say, for “unification factor”, but notice the alternatives of CNR, CCR, or CSR with $N$, $C$, and $S$ standing for “nucleus”, “core”, and “seed”.
By generalizing (2.1) we allow to use $k\times l$ CUR generators for $k$ and $l$ satisfying (1.1) and to choose any $l\times k$ nucleus $U$ for which the error matrix E=CURME=CUR-M has smaller norm.
Given two matrices $C$ and $R$, the minimal Frobenius error norm of CUR LRA
[…]
is reached for the nucleus
$U=C^{+}MR^{+}$ (2.2)
(see [36, Eqn. (6)]). We, however, cannot compute such a nucleus superfast and instead seek canonical CUR LRA (cf. [16, 40]) whose nucleus is the Moore-Penrose pseudo inverse of the $r$-truncation of a given CUR generator:
$U:=G_{r}^{+}.$ (2.3)
Given a generator $G$ we can compute $G_{r}^{+}$ by using about $kl$ scalars and $O(kl\min\{k,l\})$ flops.
We do not use the following result, but it can be of independent interest. We prove it in Appendix E.
Theorem 2.1.
[A necessary and sufficient criterion for CUR decomposition.] Let $M^{\prime}=CUR$ be a canonical CUR of $M$ for $U=G_{r}^{+}$, $G=M_{\mathcal{I},\mathcal{J}}$. Then $M^{\prime}=M$ if and only if $\operatorname{rank}(G)=\operatorname{rank}(M)$.
3 Some superfast CUR LRA algorithms
Next we present some superfast CUR LRA algorithms as heuristic. Their crude output accuracy estimates in Thm. 4.1 are greatly superseded in our numerical tests of their specified versions in Sec. 6.
3.1 Primitive and Cynical algorithms
Given five integers $k$, $l$, $m$, $n$, and $r$ satisfying (1.1) and an $m\times n$ matrix $M$, sample uniformly or fix a pair of sets $\mathcal{I}$ and $\mathcal{J}$ of $k$ row and $l$ column indexes, respectively, to define a CUR generator $G$, compute its $r$-truncation $G_{r}$, build on it a canonical CUR LRA of $M$, and call this LRA algorithm Primitive. Given the sets $\mathcal{I}$ and $\mathcal{J}$, the algorithm amounts to computing the $r$-truncation $(M_{\mathcal{I},\mathcal{J}})_{r}$ of the matrix $M_{\mathcal{I},\mathcal{J}}$ and its pseudo inverse $((M_{\mathcal{I},\mathcal{J}})_{r})^{+}$; this involves about $kl$ scalars and $O(kl\min\{k,l\})$ flops.
The following CUR LRA algorithm (we call it Cynical) [^9] Here we allude to the benefits of the austerity and simplicity of primitive life, advocated by Diogenes the Cynic, and not to shamelessness and distrust associated with modern cynicism. first fixes or samples uniformly a larger submatrix $M^{\prime}$ of $M$ and then compresses it into a $k\times l$ CUR generator of $M$ by applying one of the algorithms of [19, 41, 16].
Figure 1: A cynical CUR algorithm (the stripes mark a $p\times q$ submatrix; a $k\times l$ CUR generator is shown in black).
Algorithm 3.1.
For an $m\times n$ matrix $M$, a target rank $r$, and four integers $k$, $l$, $p$, $q$ such that
$0<r\leq k\leq p\leq m,~r\leq l\leq q\leq n,~{\rm and}~kl<pq,$ (3.1)
uniformly sample a pair of sets $\mathcal{I}$ and $\mathcal{J}$ of $p$ row and $q$ column indexes, respectively, compute a $k\times l$ CUR generator $G_{k,l}$ for the $p\times q$ submatrix $M_{\mathcal{I},\mathcal{J}}$ by applying to it one of the algorithms of [19, 41, 16], compute the $r$-truncation $G=G_{k,l,r}$, and build on it CUR LRA of $M$.
The algorithm uses about $pq$ scalars and $O(pq\min\{p,q\})$ flops and turns into the Primitive algorithm for $p=k$ and $q=l$.
3.2 Cross-Approximation (C–A) iterations
We refer to [23, 3, 40, 1], and the references therein for variation and performance of C-A iterations. In particular their maxvol version in [23] involves $O((m+n)r)$ flops per C-A iteration seeking an $r\times r$ CUR generator for an $m\times n$ matrix. Extensive effort has been invested into estimation of the overall number of C-A iterations required for convergence (see [25, 23, 40, 1] and the references therein), but adequate support of their superfast empirical performance is still a challenge. [^10] Goreinov et al proved in [23] that every C-A step of maxvol monotone increases the volume $|\det(G)|$, and reasonable bounds on the output error norms at a global maximum of the volume have been proved in [25, 40, 1]. Proving that maxvol yields global maximum superfast is still a challenge, however, although such bounds hold whp where an input matrix $M$ lies close to a factor-Gaussian matrix (see [47]).
Initialization recipes. Adaptive C-A iterations are initialized based on Gaussian elimination with complete pivoting (at a superlinear cost) combined with dynamic search for gaps in the spectrum of the singular values of $M$. Empirically the iterations converge very fast for a large and important class of inputs coming from the study of ODE and PDE (cf. [3] and the references therein). Maxvol variant of C-A iterations of [23] for $q=p=k=l=r$ is initialized by using $O(nr^{2})$ flops and still empirically converges fast for a large class of inputs, although the authors admit that their initialization recipe may not work and then suggest using costly Gaussian elimination with complete pivoting. For heuristic initialization one can alternatively try using the outputs of the algorithms of [34, 65, 15].
3.3 Maxvol for rank-1 approximation and norm estimation
For a vector ${\bf v}=(v_{i})^{n}_{i=1}\in\mathbb{R}^{n}$ and a matrix $M=(m_{i,j})_{i,j=1}^{m,n}\in\mathbb{R}^{m\times n}$ write
[…]
Let $|M|:=\max_{i,j=1}^{m,n}|m_{i,j}|$ define the inconsistent matrix norm [^11] E.g., $|M^{2}|=n>|M|^{2}=1$ for $M=(1)_{i,j=1}^{n,n}$. $|M|:=\max_{i,j=1}^{m,n}|m_{i,j}|$, called Chebyshev’s norm in [40], such that
$|M|\leq||M||\leq\sqrt{mn}~|M|,~|M|\leq|||M||_{1}\leq m|M|,~|M|\leq||M||_{\infty}\leq n|M|.$ (3.3)
One can compute the norms $|M|$, $_{1}, M||M||_{\infty}$ by using 3mn33mn-3 additions and comparisons.
For $p=q=r=1$ maxvol computes an absolutely largest entry of $M=(m_{i,j})_{i,j=1}^{m,n}$ in its fixed column or row (in its “vertical” or “horizontal” step), respectively, by using m1m-1 or n1n-1 comparisons, respectively. Maxvol stops at the entry $m_{i,j}$ that maximizes both $|m_{g,j}|$ over all $g$ and $|m_{i,h}|$ over all $h$. For a large class of matrices $M$ this is also the global maximum $|m_{i,j}|=|M|$.
In that case Eqns. (3.2) and (3.3) imply that the 1-norms of the $j$th column vector $C:=(m_{g,j})_{g=1}^{m}$ and the $i$th row vector $R:=(m_{i,h})_{h=1}^{n}$ provide crude estimates for the 1-norm and the $\infty$-norm of $M$, respectively, being off by at most factors of $m$ and $n$, respectively.
Furthermore, write $U:=1/m_{i,j}$, obtain rank-1 approximation of $M$ with the matrix $CUR$, and recall from the result of [1], extending [25, 40], that
$|M-CUR|\leq\frac{2\sigma_{1}(M)\sigma_{2}(M)}{(\sigma_{1}^{2}(M)+\sigma^{2}_{2}(M))^{1/2}}.$
Thm. 1.1 and bound (3.3) combined imply that this approximation is optimal under the spectral norm up to a factor of $2\sqrt{mn}$.
4 A posteriori errors of a canonical CUR LRA
Example A.1 shows that a posteriori errors of an LRA given by a matrix $M^{\prime}\approx M$ cannot be estimated superfast for worst case inputs, but next we prove such a bound where $M^{\prime}$ is given by a triplet $\{C,U,R\}$.
4.1 Error Estimation: an Outline and the Statement
We estimate the error norm of CUR rank-$r$ approximation of $M$ by comparing it with CUR decomposition of a nearby rank-$r$ matrix $M^{\prime}$ defined by the same sets of row and column indexes.
Outline 4.1.
[Error Estimation for a Canonical CUR LRA.]
  1. 1.
    Consider (but do not compute) an auxiliary $m\times n$ matrix $M^{\prime}$ of rank $r$ that approximates the matrix $M$ within a fixed norm bound $\epsilon$ such that
[…]
    [E.g., we can choose $M^{\prime}=M_{r}$ and $\epsilon:=\sigma_{r+1}(M)$.]
  2. 2.
    For the matrices $M$ and $M^{\prime}$ fix two row and column index sets $\mathcal{I}$ and $\mathcal{J}$, respectively, and define $k\times l$ generators $G=M_{\mathcal{I},\mathcal{J}}$ and $G^{\prime}=M_{\mathcal{I},\mathcal{J}}^{\prime}$, nuclei $U=G_{r}^{+}$ and $U^{\prime}=G_{r}^{\prime+}$, and canonical CUR approximation $M\approx CUR$ and decomposition $M^{\prime}=C^{\prime}U^{\prime}R^{\prime}$.
  3. 3.
    Observe that
[…]
  4. 4.
    Bound the norm $C^{\prime}U^{\prime}R^{\prime}-CUR||$ in terms of the values $\epsilon$, $C||$, $U||$, and $R||$.
Next we elaborate upon step 4 provided that we have already performed steps 1 – 3.
Theorem 4.1.
Given an $m\times n$ matrix $M$, a $k\times l$ matrix $G:=M_{\mathcal{I},\mathcal{J}}$, a positive $\epsilon\leq\sigma_{r}(G)$, a positive integer $r<\min\{k,l\}$, and an $m\times n$ rank-$r$ matrix $M^{\prime}$ satisfying (4.1), fix $v$ of (1.2), write
$C:=M_{:,\mathcal{J}},~R:=M_{\mathcal{I},:},~U=G_{r}^{+},$ (4.3)
$\zeta:=\begin{cases}\sqrt{2}&{\rm for}~~r=\min\{k,l\}\\ (1+\sqrt{5})/2&{\rm for}~~r<\min\{k,l\}\end{cases}$ (4.4)
[…]
$M-CUR||\leq(v+1)\Big{(}\frac{2\zeta}{1-\theta}~(v+1)+2\Big{)}\epsilon,~{\rm where}~2\zeta\leq 1+\sqrt{5}.$ (4.5)
In view of (4.2) we only need to estimate the norm $C^{\prime}U^{\prime}R^{\prime}-CUR||$.
4.2 The first bound on $C^{\prime}U^{\prime}R^{\prime}-CUR||$
Lemma 4.1.
Fix five integers $k$, $l$, $m$, $n$, and $r$ such that $r\leq k\leq m$ and $r\leq l\leq n$, an $m\times n$ matrix $M$, its rank-$r$ approximation $M^{\prime}$ satisfying (4.1), and canonical CUR LRAs
$M\approx CUR~{\rm and}~M^{\prime}=C^{\prime}U^{\prime}R^{\prime}$
defined by the same pair of index sets $\mathcal{I}$ and $\mathcal{J}$ of cardinality $k$ and $l$, respectively, such that
$C:=M_{:,\mathcal{J}},~R:=M_{\mathcal{I},:},~U=G_{r}^{+},~~C^{\prime}:=M^{\prime}_{:,\mathcal{J}},~R^{\prime}:=M^{\prime}_{\mathcal{I},:},~U^{\prime}=G_{r}^{\prime+},$
$G=M_{\mathcal{I},\mathcal{J}},~{\rm and}~G^{\prime}=M^{\prime}_{\mathcal{I},\mathcal{J}}.$
Then
[…]
Proof.
Notice that
$CUR-C^{\prime}U^{\prime}R^{\prime}=(C-C^{\prime})UR+C^{\prime}U(R-R^{\prime})+C^{\prime}(U-U^{\prime})R^{\prime}.$
Therefore
[…]
4.3 Estimation of the norm $U-U^{\prime}||$
Lemma 4.2.
Under the assumptions of Lemma 4.1 we have
[…]
Proof.
Recall that $\operatorname{rank}(G^{\prime})=\operatorname{rank}(M^{\prime})=\operatorname{rank}(G_{r})=r$ and that
[…]
Then apply [4, Thm. 2.2.5] for $A=G,~B=G^{\prime}$. ∎
Lemma 4.3.
[…]
Proof.
Thm. 1.2 implies that
[…]
Substitute (1.2) and obtain that $\sigma_{r}(G^{\prime})\geq(1-\theta)\sigma_{r}(G)>0$. Hence
[…]
Corollary 4.1.
[…]
4.4 Proof of Thm. 4.1
Combine bound (4.2) and Lemma 4.1 and obtain
[…]
Combine (1.2) and Cor. 4.1 and deduce that
$||C^{\prime}||~||U-U^{\prime}||\leq(||C||+\epsilon)\frac{2\zeta}{1-\theta}~||U||^{2}\epsilon<\frac{2\zeta}{1-\theta}(||C||~|U||^{2}+||U||)\epsilon.$
Hence
[…]
Substitute relationships (4.3) and obtain
[…]
Combine this bound with Eqns. (4.6) and (4.7) and obtain the theorem.
5 The errors of CUR LRA of a perturbed factor-Gaussian matrix
In this section we prove that a canonical CUR LRA of a matrix $M$ with a generator $G=M_{\mathcal{I},\mathcal{J}}$ is quite accurate whp for any fixed pair of sufficiently large sets $\mathcal{I}$ and $\mathcal{J}$ of rows and columns indexes if $M$ is close to a two-sided factor-Gaussian matrix $M^{\prime}$ of low rank. Unlike Def. 1.1 we use character $H$ for Gaussian matrices, leaving character $G$ for CUR generators.
Theorem 5.1.
Given two independent Gaussian matrices $H_{1}\in\mathbb{R}^{m\times\rho}$ and $H_{2}\in\mathbb{R}^{\rho\times n}$ and a $\rho\times\rho$ (well-conditioned) matrix $\Sigma$ of full rank $\rho$ such that
$k<m,~l<n,~\textrm{and}~\min\{k,l\}\geq\rho+2\geq 4,$ (5.1)
write
$M^{\prime}:=H_{1}\Sigma H_{2},$ (5.2)
$C^{\prime}:=M^{\prime}_{:,\mathcal{J}},~R^{\prime}:=M^{\prime}_{\mathcal{I},:},~G^{\prime}=M^{\prime}_{\mathcal{I},\mathcal{J}}\in\mathbb{R}^{k\times l},~{\rm and}~U^{\prime}:=G^{\prime+}$
for fixed row and column index sets $\mathcal{I}$ and $\mathcal{J}$. Then $C^{\prime}$, $R^{\prime}$, and $G^{\prime}$ are factor-Gaussian matrices such that
[…]
for  $e:=2.7182818\dots$,
$\alpha:=\frac{e^{2}\sqrt{kl}}{(k-\rho)(l-\rho)}~\textrm{and}~\beta:=\max\big{\{}(\sqrt{m}+\sqrt{\rho})(\sqrt{\rho}+\sqrt{l}),(\sqrt{n}+\sqrt{\rho})(\sqrt{\rho}+\sqrt{k})\big{\}}.$ (5.3)
Proof.
[…]
Recall that $C^{\prime}=M^{\prime}_{:,\mathcal{J}}=H_{1}\Sigma{H_{2}}_{:,\mathcal{J}}$ and $H_{1}$ and $H_{2}$ are independent of one another and deduce that
[…]
Recall from Thm. C.1 that $\mathbb{E}(\nu_{p,q})\leq\sqrt{p}+\sqrt{q}$ and obtain
[…]
Next estimate $U^{\prime}||$. Apply Lemma 1.1, recall from Thm. C.2 that
$\mathbb{E}(\nu^{+}_{p,q})\leq\frac{e\sqrt{p}}{p-q},~{\rm for}~p\geq q+2\geq 2,$
and deduce that
[…]
$\leq\frac{e^{2}\sqrt{kl}}{(k-\rho)(l-\rho)\sigma_{\rho}(\Sigma)}=\frac{\alpha}{\sigma_{\rho}(\Sigma)}.$
Now assume that an input matrix $M$ is a perturbed two-sided factor-Gaussian matrix and that two fixed sets $\mathcal{I}$ and $\mathcal{J}$ of its row and column indexes are sufficiently large and then estimate the norms $U||$, $C||$ and $R||$.
Lemma 5.1.
Under the assumptions of Thm. 5.1, let $\epsilon\leq\frac{\sigma_{\rho}(\Sigma)}{30\alpha}$ be a small positive number and let $E\in\mathbb{R}^{m\times n}$ be a perturbation matrix such that $E||\leq\epsilon$. Write
$M:=M^{\prime}+E,~C:=M_{:,\mathcal{J}},~R:=M_{\mathcal{I},;},~{\rm and}~U:=(M_{\mathcal{I},\mathcal{J}})_{\rho}^{+}.$
Then with a probability no less than 0.7 we have
[…]
Proof.
[…]
$U^{\prime}||\leq\frac{10\alpha}{\sigma_{\rho}(\Sigma)}\textrm{ or equivalently }\sigma_{\rho}(M^{\prime}_{\mathcal{I},\mathcal{J}})\geq\frac{\sigma_{\rho}(\Sigma)}{10\alpha}$ (5.4)
with a probability no less than 0.9.
Similarly to the argument of Lemma 4.3, deduce that
$\sigma_{\rho}(M_{\mathcal{I},\mathcal{J}})\geq\sigma_{\rho}(M^{\prime}_{\mathcal{I},\mathcal{J}})-\epsilon\geq\frac{\sigma_{\rho}(\Sigma)}{15\alpha},$
and then claim (i) follows.
[…]
with a probability no less than 0.8. Then combine claim (i) with the union bound and obtain that
[…]
with a probability no less than 0.7. ∎
Now choose any generator of CUR LRA of a perturbed factor-Gaussian matrix of rank $\rho$ and prove a probabilistic upper bound on the error of a resulting CUR LRA of that matrix.
Theorem 5.2.
If the assumptions of Lemma 5.1 hold for a perturbed factor-Gaussian matrix $M$, then bound (4.5) on the approximation error norm MCUR||M-CUR|| for $\theta\leq 1/2$, that is,
$M-CUR||\leq(4\zeta v+2)(v+1)\epsilon,~{\rm for}~4\zeta\leq 2+2\sqrt{5},~v=150\alpha\beta\sigma_{1}(\Sigma)/\sigma_{\rho}(\Sigma)+1/2,$ (5.5)
holds with a probability no less than 0.6.
Proof.
Combine bounds $\epsilon\leq\frac{\sigma_{\rho}(\Sigma)}{30\alpha}$ and (5.4) of Lemma 5.1 and obtain that $\sigma_{\rho}(G)\geq 3\epsilon>\epsilon$ and therefore (4.5) hold with a probability $p_{0}\geq 0.9$. Replace $\theta$ and $v$ with the upper bounds 1/2 and $150\alpha\beta\sigma_{1}(\Sigma)/\sigma_{\rho}(\Sigma)+1/2$, respectively, holding with a probability $p_{1}\geq 0.7$. Combine this bound with $p_{0}\geq 0.9$. ∎
Remark 5.1.
Recall that
$\alpha:=\frac{e^{2}\sqrt{kl}}{(k-\rho)(l-\rho)}~{\rm and}~\beta:=\max\big{\{}(\sqrt{m}+\sqrt{\rho})(\sqrt{\rho}+\sqrt{l}),(\sqrt{n}+\sqrt{\rho})(\sqrt{\rho}+\sqrt{k})\big{\}}.$
Furthermore, suppose that $m\gg k\gg\rho$, $n\gg l\gg\rho$, and
$v=150\alpha\beta\sigma_{1}(\Sigma)/\sigma_{\rho}(\Sigma)+1/2=O\Big{(}\max\Big{\{}\frac{\sqrt{m}}{\sqrt{k}},\frac{\sqrt{n}}{\sqrt{l}}\Big{\}}\Big{)}.$
Now drop the smaller terms of (5.5), and obtain that its dominant part is
$O\Big{(}\max\Big{\{}\frac{m}{k},\frac{n}{l}\Big{\}}\cdot\epsilon\Big{)}.$
6 Numerical experiments
6.1 Test overview
For both synthetic and real world input matrices the test results in our Table 6.1 show good accuracy of CUR LRA output by just five C-A loops (with uniform initialization). Accuracy was still quite good where we applied the Cynical algorithm but significantly worse where we applied the Primitive algorithm. Tables 6.2 and 6.3 compare the output accuracy of the random sampling algorithms of [16] and their superfast combination with C-A iterations.
We have performed the tests in the Graduate Center of the City University of New York by applying MATLAB. We generated synthetic inputs by using Gaussian matrices, which we created by using standard normal distribution function “randn()”. We calculated $\epsilon$-ranks of matrices for $\epsilon=10^{-6}$ by using the MATLAB’s function ”rank(-,1e-6)”, which only counts singular values greater than $10^{-6}$. Some numerical experiments were executed with software custom programmed in $C^{++}$ and compiled with LAPACK version 3.6.0 libraries.
Our tables display the mean value of the relative output errors $\frac{\|M-\tilde{M}\|}{\|M\|}$ for an input matrix $M$ and output $\tilde{M}:=CUR$ and the standard deviation (std) over 1000 runs for every class of inputs.
6.2 Input matrices for LRA
We used the following two classes of input matrices $M$ for testing LRA algorithms.
Class I (Synthetic inputs): $n\times n$ random matrices $M$ having expected rank $r$, that is,
$M=G_{1}*G_{2}+10^{-10}G_{3},$
for three Gaussian matrices $G_{1}$ of size $n\times r$, $G_{2}$ of size $r\times n$, and $G_{3}$ of size $n\times n$.
[…]
namely, to the following six input families from the Database:
baart: Fredholm Integral Equation of the first kind,
shaw: one-dimensional image restoration model,
gravity: 1-D gravity surveying model problem,
wing: problem with a discontinuous solution,
foxgood: severely ill-posed problem,
inverse Laplace: inverse Laplace transformation.
6.3 Four test algorithms
We tested the following four algorithms for computing CUR LRA of input matrices $M$ having $\epsilon$-rank $r$:
  • Tests 1 (The Primitive algorithm for $k=l=r$): Uniformly sample two index sets $\mathcal{I}$ and $\mathcal{J}$, both of cardinality $r$, then compute a nucleus $U=M_{\mathcal{I},\mathcal{J}}^{-1}$ and define CUR LRA
    $\tilde{M}:=CUR=M_{:,\mathcal{J}}\cdot M_{\mathcal{I},\mathcal{J}}^{-1}\cdot M_{\mathcal{I},\cdot}.$ (6.1)
  • Tests 2 (Five loops of C–A): Uniformly sample an initial row index set $\mathcal{I}_{0}$ of cardinality $r$, then perform five loops of C–A (cf. Sec. 3.2) incorporating Alg. 1 of [41] as a subalgorithm that produces $r\times r$ CUR generators. At the end compute a nucleus $U$ and define CUR LRA as in Tests 1.
  • Tests 3 (A Cynical algorithm for $p=q=4r$ and $k=l=r$): Uniformly sample a row index set $\mathcal{K}$ and a column index set $\mathcal{L}$, both of cardinality $4r$. Then apply Algs. 1 and 2 of [41] to compute an $r\times r$ submatrix $M_{\mathcal{I},\mathcal{J}}$ of $M_{\mathcal{K},\mathcal{L}}$. Compute a nucleus and obtain CUR LRA of (6.1).
  • Tests 4 (Combination of a single C–A loop with Tests 3): Uniformly sample a column index set $\mathcal{L}$ of cardinality $4r$; then perform a single C–A loop: first define an index set $\mathcal{K}^{\prime}$ of cardinality $4r$ and the submatrix $M_{\mathcal{K}^{\prime},\mathcal{L}}$ in $M_{:,\mathcal{L}}$ by applying Alg. 1 of [41]; then apply this algorithm to the matrix $M_{\mathcal{K}^{\prime},:}$ to find an index set $\mathcal{L}^{\prime}$ of cardinality $4r$ and to define submatrix $M_{\mathcal{K}^{\prime},\mathcal{L}^{\prime}}$ in $M_{\mathcal{K}^{\prime},:}$. Then proceed as in Tests 3 – find an $r\times r$ CUR generator $M_{\mathcal{I},\mathcal{J}}$ in $M_{\mathcal{K}^{\prime},\mathcal{L}^{\prime}}$ by applying Algs. 1 and 2 of [41] and finally compute a nucleus and CUR LRA of $M$ based on (2.3) for $G:=M_{\mathcal{I},\mathcal{J}}$.
6.4 CUR LRA of the matrices of class I
Table 6.1 shows the summary statistics for each of the four tests matrices of class I for $n=256,512,1024$ and $r=8,16,32$. In the column SVD we display the data for the best rank-$r$ approximation given by the $(r+1)$-st largest singular value to establish the baseline for the performance of our tests. The results of Tests 1 fall in the range $10^{-8},10^{-7}]$, those of Tests 2, 3 and 4 in $10^{-11},10^{-10}]$ in the decreasing order of accuracy. On the average the error bounds of Tests 2, 4, 3, and 1 exceed the baseline bounds in the column SVD by roughly factors of 10, 15, 30, and over 10,000, respectively, but even the crudest LRAs output in Test 1 could be used in some applications.
SVD Tests 1 Tests 2 Tests 3 Tests 4
n r mean std mean std mean std mean std mean std
256 8 1.01e-11 3.92e-13 1.60e-08 8.65e-08 5.94e-11 8.06e-12 1.13e-10 2.36e-11 8.23e-11 1.64e-11
256 16 9.12e-12 3.09e-13 2.44e-07 5.85e-06 7.31e-11 1.02e-11 1.12e-10 2.02e-11 9.45e-11 1.51e-11
256 32 7.80e-12 2.32e-13 4.82e-08 2.50e-07 8.93e-11 1.10e-11 1.13e-10 1.69e-11 1.04e-10 1.50e-11
512 8 7.64e-12 2.23e-13 3.50e-08 2.72e-07 5.71e-11 7.08e-12 1.21e-10 2.67e-11 8.34e-11 1.62e-11
512 16 7.06e-12 1.68e-13 1.18e-07 2.53e-06 7.08e-11 8.96e-12 1.26e-10 2.21e-11 9.98e-11 1.59e-11
512 32 6.36e-12 1.37e-13 7.43e-08 7.47e-07 9.25e-11 1.14e-11 1.34e-10 1.92e-11 1.20e-10 1.73e-11
1024 8 5.63e-12 1.13e-13 2.42e-08 2.46e-07 5.39e-11 5.89e-12 1.28e-10 2.83e-11 8.10e-11 1.55e-11
1024 16 5.34e-12 9.23e-14 6.12e-08 6.91e-07 6.94e-11 7.68e-12 1.37e-10 2.35e-11 1.04e-10 1.73e-11
1024 32 4.95e-12 7.55e-14 6.20e-07 1.36e-05 9.17e-11 1.06e-11 1.51e-10 2.09e-11 1.29e-10 1.87e-11
Table 6.1: Errors of CUR LRA of random matrices of class I
6.5 LRA by means of random sampling and C-A acceleration
Tables 6.2 and 6.3 display the relative errors $\frac{\|M-\tilde{M}\|}{\|M\|}$ of the LRA $\tilde{M}$ of the matrices $M$ computed in two ways: by means of [16, Alg. 2] (see the lines marked “CUR”), which are expected to be near-optimal under the Frobenius morm whp, or eight C-A iterations combined with [16, Alg. 1] (see the lines marked “C-A”). The algorithms run at superlinear cost in the former case and are superfast in the latter case. The columns of the tables marked with $\epsilon$-rank” display $\epsilon$-rank of an input matrix. The columns of the tables marked with ”k=lk=l” show the number of rows and columns in a square matrix of CUR generator.
As we pointed out in Sec. 1.2.1, in almost all our tests this dramatic acceleration was achieved at the price of only minor deterioration of output accuracy.
input algorithm m n $\epsilon$-rank k=l mean std
finite diff C-A 608 1200 94 376 6.74e-05 2.16e-05
finite diff CUR 608 1200 94 376 6.68e-05 2.27e-05
finite diff C-A 608 1200 94 188 1.42e-02 6.03e-02
finite diff CUR 608 1200 94 188 1.95e-03 5.07e-03
baart C-A 1000 1000 6 24 2.17e-03 6.46e-04
baart CUR 1000 1000 6 24 1.98e-03 5.88e-04
baart C-A 1000 1000 6 12 2.05e-03 1.71e-03
baart CUR 1000 1000 6 12 1.26e-03 8.31e-04
baart C-A 1000 1000 6 6 6.69e-05 2.72e-04
baart CUR 1000 1000 6 6 9.33e-06 1.85e-05
shaw C-A 1000 1000 12 48 7.16e-05 5.42e-05
shaw CUR 1000 1000 12 48 5.73e-05 2.09e-05
shaw C-A 1000 1000 12 24 6.11e-04 7.29e-04
shaw CUR 1000 1000 12 24 2.62e-04 3.21e-04
shaw C-A 1000 1000 12 12 6.13e-03 3.72e-02
shaw CUR 1000 1000 12 12 2.22e-04 3.96e-04
Table 6.2: LRA errors of Cross-Approximation (C-A) incorporating [16, Algorithm 1] in comparison to stand-alone CUR of [16, Algorithm 2].
input algorithm m = n $\epsilon$-rank $k=l$ mean std
foxgood C-A 1000 10 40 3.05e-04 2.21e-04
foxgood CUR 1000 10 40 2.39e-04 1.92e-04
foxgood C-A 1000 10 20 1.11e-02 4.28e-02
foxgood CUR 1000 10 20 1.87e-04 4.62e-04
wing C-A 1000 4 16 3.51e-04 7.76e-04
wing CUR 1000 4 16 2.47e-04 6.12e-04
wing C-A 1000 4 8 8.17e-04 1.82e-03
wing CUR 1000 4 8 2.43e-04 6.94e-04
wing C-A 1000 4 4 5.81e-05 1.28e-04
wing CUR 1000 4 4 1.48e-05 1.40e-05
gravity C-A 1000 25 100 1.14e-04 3.68e-05
gravity CUR 1000 25 100 1.41e-04 4.07e-05
gravity C-A 1000 25 50 7.86e-04 4.97e-03
gravity CUR 1000 25 50 2.22e-04 1.28e-04
inverse Laplace C-A 1000 25 100 4.15e-04 1.91e-03
inverse Laplace CUR 1000 25 100 5.54e-05 2.68e-05
inverse Laplace C-A 1000 25 50 3.67e-01 2.67e+$0$
inverse Laplace CUR 1000 25 50 2.35e-02 1.71e-01
Table 6.3: LRA errors of Cross-Approximation (C-A) incorporating [16, Algorithm 1] in comparison to stand-alone CUR of [16, Algorithm 2].
PART II: LRA VIA RANDOM SKETCHING
7 LRA by means of sketching: four algorithms
Next we slightly generalize the sketching LRA algorithms of [30, 60] (see Remarks 7.17.3); Thms. 8.2 and 9.1 combined formally support output accuracy of dual version of all of them whp; we applied our tests to the algorithms of [30, 60] accelerated by means of using Abridged SRHT multipliers $F$ and $H$.
Algorithm 7.1.
Input:
An $m\times n$ matrix $M$ and a target rank $r$.
Output:
Two matrices $X\in\mathbb{R}^{m\times l}$ and $Y\in\mathbb{R}^{l\times n}$ for $r\leq l\leq n$ defining an LRA $\tilde{M}=XY$ of $M$.
Initialization:
Fix an integer $l=r+p\leq n$, for $p\geq 0$, and an $n\times l$ matrix $H$ of full rank $l$.
Computations:
  1. 1.
    Compute the $m\times l$ matrix $MH$.
  2. 2.
    Fix a nonsingular $l\times l$ matrix $T^{-1}$ and output the $m\times l$ matrix $X:=MHT^{-1}$.
  3. 3.
[…]
Algorithm 7.2.
Input:
As in Alg. 7.1.
Output:
Two matrices $X\in\mathbb{R}^{k\times n}$ and $Y\in\mathbb{R}^{m\times k}$ defining an LRA $\tilde{M}=YX$ of $M$.
Initialization:
Fix an integer $k=r+p\leq m$, for $p\geq 0,$ and a $k\times m$ matrix $F$ of full numerical rank $k$.
Computations:
  1. 1.
    Compute the $k\times m$ matrix $FM$.
  2. 2.
    Fix a nonsingular $k\times k$ matrix $S^{-1}$; then output $k\times n$ matrix $X:=S^{-1}FM$.
  3. 3.
[…]
The following algorithm combines row and column sketching.
Algorithm 7.3.
Input:
As in Alg. 7.1.
Output:
Two matrices $X\in\mathbb{R}^{m\times k}$ and $Y\in\mathbb{R}^{k\times n}$ defining an LRA $\tilde{M}=XY$ of $M$.
Initialization:
Fix two integers $k$ and $l$, $r\leq k\leq m$ and $r\leq l\leq n$; fix two matrices $F\in\mathbb{R}^{k\times m}$ and $H\in\mathbb{R}^{n\times l}$ of full numerical ranks and two nonsingular matrices $S\in\mathbb{R}^{k\times k}$ and $T\in\mathbb{R}^{l\times l}$.
Computations:
1. Output the matrix $X=MHT^{-1}\in\mathbb{R}^{m\times l}$.
2. Compute the matrices $U:=S^{-1}FM\in\mathbb{R}^{k\times n}$ and $W:=S^{-1}FX\in\mathbb{R}^{m\times l}$.
[…]
Remark 7.1.
We can obtain Algs. 7.2 by applying Alg. 7.1 to the transpose $M^{*}$. Likewise, by applying Alg. 7.3 to $M^{*}$ we obtain Alg. 6.4. We only study Algs. 7.1 and 7.3, but can readily extend that study to Algs. 7.2 and 6.4.
Remark 7.2.
Fix $k=r$, $l>k$, a random $n\times l$ matrix $H$ (e.g., Gaussian or SRHT matrix), identity matrix $S$, and $T$ equal to the $R$ factor in the $QR$ factorization of $MH$. Then the matrix $X$ has orthonormal columns and Algs. 7.1 turns into Proto-algorithm of [30, Sec. 9], while Alg. 7.3 turns into the Generalized Nyström algorithm of [39, Eqn. (3)]. [39] stabilizes the latter algorithm numerically – essentially by means of setting to 0 all singular values of the matrix $W$ exceeded by a fixed $\epsilon$, “a modest multiple of the unit roundoff $u$ times W||W||”. Our study can be readily extended to such a stabilized LRA because the stabilization little affects complexity of LRA and only improves its output accuracy. We can obtain various other modifications of Algs. 7.1 and 7.3 by fixing other sketching matrices $F$ and $H$.
Remark 7.3.
Column (resp. row) sketching turns into column (resp. row) subset selection where $H$ in Alg. 7.1 (resp. $F$ in Alg. 7.2) is a sampling matrix, that is, a full rank submatrix of a permutation matrix.
8 Deterministic output error bounds for sketching algorithms
8.1 Deterministic error bounds of Range Finder
Next we recall some known estimates for the errors of Alg. 7.1, to be used in the next section.
Theorem 8.1.
[30, Thm. 9.1]. Suppose that Alg. 7.1 has been applied to a matrix $M$ and let
$M=\begin{pmatrix}U_{1}&U_{2}\end{pmatrix}\begin{pmatrix}\Sigma_{1}&\\ &\Sigma_{2}\end{pmatrix}\begin{pmatrix}~V_{1}^{*}\\ ~V_{2}^{*}\end{pmatrix}~~{\rm and}~M_{r}=U_{1}\Sigma_{1}V_{1}^{*}$
be SVDs of the matrices $M$ and its rank-$r$ truncation $M_{r}$, respectively. $\Sigma_{2}=O$ and $XY=M$ if $\operatorname{rank}(M)=r$. The $r$ columns of $V_{1}$ are the $r$ top right singular vectors of $M$.] Write
$C_{1}=V^{*}_{1}H,~C_{2}=V^{*}_{2}H.$ (8.1)
Assume that $H||_{2}\leq 1$ and $\operatorname{rank}(C_{1})=r$. Then
[…]
Corollary 8.1.
Under the assumptions of Thm. 8.1 and for $\tau_{r+1}(M)$ of Thm. 1.1 it holds that
[…]
Proof.
The corollary follows from (8.2) because
[…]
8.2 Impact of pre-multiplication on the errors of LRA
The following theorem shows that the overall error bounds of Alg. 7.3 are dominated by the product of the norm $|W|||$ and the error norm bound of Alg. 7.1.
Theorem 8.2.
See [60]. Let Alg. 7.3 output a matrix $XY$ for $Y=(FX)^{+}FM$ and let $m\geq k\geq l=\operatorname{rank}(X)$. Then
$M-XY=W(M-XX^{+}M)~{\rm for}~W=I_{m}-X(FX)^{+}F,$ (8.4)
[…]
Proof.
Recall that $Y=(FX)^{+}FM$ and notice that $(FX)^{+}FX=I_{l}$ if $k\geq l=\operatorname{rank}(FX)$. Therefore, $Y=X^{+}M+(FX)^{+}F(M-XX^{+}M)$. Consequently, (8.4) and (8.5) hold. ∎
Remark 8.1.
[…]
if both matrices $F$ and $X$ have orthonormal columns. The sketch matrix $F$ is our choice, and for $X$ we ensure column orthogonality by properly choosing the matrix $T$ in Alg. 7.3 (see Remark 7.2).
9 Output error norm bounds for dual sketching algorithms
Given the matrices $MHT^{-1}$ and $S^{-1}FM$, Alg. 7.3 uses $O(kln)$ flops and hence runs superfast for $kl\ll m$. If also $l^{2}\ll m$ and $k^{2}\ll n$, then for proper Ultrasparse matrices $F$ and $H$ we can compute the matrices $MHT^{-1}$ and $S^{-1}FM$ superfast as well and hence can perform entire Algorithm 7.3 superfast, involving much less than $mn$ entries of $M$ and other scalars, already where $k\ll m$ and $l\ll n$. Although such an LRA fails for a worst case input, it succeeds whp under our dual model in the case of any fixed, possibly sparse, well-conditioned multipliers $F$ and $H$ of full rank and a random input matrix $M$ that admits LRA. In view of the previous subsection we only need to prove this for Alg. 7.1.
9.1 Auxiliary results
Lemma 9.1.
Suppose that $G$ and $H$ are $r\times n$ and $n\times l$ matrices, respectively, $r<l<n$, GHGH has full rank $r$, and $Q$ is an $r\times n$ matrix with orthonormal rows such that $Q$ and $G$ have the same row space. Then
$\sigma_{r}(QH)\geq\frac{\sigma_{r}(GH)}{\sigma_{1}(G)}.$
Proof.
Without loss of generality, suppose that $R\in\mathbb{R}^{r\times r}$ and $G=RQ$. Then
$\sigma_{r}(QH)=\sigma_{r}(R^{-1}GH)\geq\sigma_{r}(GH)\cdot\sigma_{r}(R^{-1}).$
Hence $\sigma_{r}(R^{-1})=1/\sigma_{1}(G)$ because the matrices $R$ and $G$ share their singular values. ∎
Lemma 9.2.
Suppose that $G$ is an $r\times n$ Gaussian matrix, $H$ is an $n\times l$ matrix with orthonormal columns, $n>l>r$, $l\geq 4$, $Q$ is a matrix with orthonormal rows, and $Q$ and $G$ share their row space. Fix two positive parameters $t_{1}$ and $t_{2}<1$. Then
$\sigma_{r}(QH)\geq t_{2}\cdot\frac{\sqrt{l}-\sqrt{r/l}+\sqrt{1/l}}{t_{1}+\sqrt{r}+\sqrt{n}}\cdot\frac{1}{e}\$
with a probability no less than $1-\exp(-t_{1}^{2}/2)-(t_{2})^{l-r}$.
Proof.
The matrix $GH$ has distribution of an $r\times l$ Gaussian matrix by virtue of Lemma 1.3, and hence we can assume that it has full rank (see Remark 1.2). Now recall Thm. C.1 and claim (ii) of Thm. C.2 and obtain
$\operatorname{Probability}\{\sigma_{1}(G)>t_{1}+\sqrt{r}+\sqrt{n}\}<\exp(-t_{1}^{2}/2)~{\rm for}~t_{1}\geq 0~{\rm and}$
$\operatorname{Probability}\Big{\{}\sigma_{r}(GH)\leq t_{2}\cdot\frac{l-r+1}{e\sqrt{l}}\Big{\}}\leq(t_{2})^{l-r}~{\rm for}~t_{2}<1~{\rm and}~l\geq 4.$
Combine the latter two inequalities, the union bound, and Lemma 9.1 and obtain Lemma 9.2. ∎
Lemma 9.2 implies that $\sigma_{r}(QH)$ has at least order of $\sqrt{l/n}$ whp.
Corollary 9.1.
For $n,l,r,G,Q$, and $H$ of Lemma 9.2, let $n>36r$ and l>22(r1)l>22(r-1). Then
$\operatorname{Probability}\big{\{}\sigma_{r}(QH)\geq\frac{1}{4}\sqrt{l/n}\big{\}}>1-\exp\Big{(}-\frac{n}{72}\Big{)}-\exp\Big{(}-\frac{l-r}{20}\Big{)}.$
Proof.
Write $t_{1}:=\frac{1}{3}\sqrt{n}-\sqrt{r}$ and $t_{2}:=\frac{1}{3}~\frac{le}{l-r+1}$, recall that $n>36r$ and l>22(r1)l>22(r-1), and then readily verify that $t_{1}>\frac{\sqrt{n}}{6}$ and $\exp(0.05)>0.95>t_{2}>0$. Finally apply Lemma 9.2 under these bounds on $t_{1}$ and $t_{2}$. ∎
Remark 9.1.
We can extend our lower bounds of Lemma 9.2 and Cor. 9.1 on $\sigma_{r}(QH)$ to the case of any matrix $H$ of full rank $l$ if we decrease these bounds by a factor of $\sigma_{l}(H)$.
9.2 Output errors of Alg. 7.1 for a matrix with random singular space
Assumption 9.1.
Let $r<n\leq m$ (we can readily extend our study to the case where $m<n$). Fix two constant matrices
$\Sigma_{r}=\operatorname{diag}(\sigma_{j})_{j=1}^{r}~{\rm and}~\Sigma_{\perp}=\begin{bmatrix}\operatorname{diag}(\sigma_{j})_{j=r+1}^{n}\\ O_{m-n,n-r}\end{bmatrix}\in\mathbb{R}^{(m-r)\times(n-r)}$
such that $\sigma_{1}\geq\sigma_{2}\geq\dots\geq\sigma_{n}\geq 0$, and $\sigma_{r}>0$. Let $G$ be an $r\times n$ Gaussian matrix and let $Q\in\mathbb{R}^{n\times r}$ and $Q_{\perp}\in\mathbb{R}^{(n-r)\times r}$ be two matrices whose column sets make up orthonormal bases of the row space of $G$ and its orthogonal complement, respectively. Furthermore, let
$M=U\cdot\begin{bmatrix}\Sigma_{r}&0\\ 0&\Sigma_{\perp}\end{bmatrix}\cdot\begin{bmatrix}Q^{*}\\ Q_{\perp}^{*}\end{bmatrix}$ (9.1)
be SVD where the right singular space of $M$ is random and $U$ is a matrix with orthonormal columns.
Remark 9.2.
The matrix $G$ does not uniquely define the matrices $Q$, $Q_{\perp}$, and $U$ under Assumption 9.1 and hence does not uniquely define the matrix $M$, but this is immaterial for our analysis.
Theorem 9.1.
[Errors of Alg. 7.1 for an input with a random singular space.] Suppose that $G$ is an $r\times n$ Gaussian matrix, $H\in\mathbb{R}^{n\times l}$ is a constant matrix, $n>36r$, l>22(r1)l>22(r-1), $r\leq l<\min(m,n)$, and Alg. 7.1 applied to the matrix $M$ of (9.1) outputs two matrices $X$ and $Y$.
(i) If $H$ has orthonormal columns, then
[…]
with a probability no less than $1-\exp(-\frac{n}{72})-\exp(-\frac{l-r}{20})$.
(ii) If $H$ has full rank $l$, then
[…]
with a probability no less than $1-\exp(-\frac{n}{72})-\exp(-\frac{l-r}{20})$.
Proof.
We can assume that the matrices $G$ and $GH$ have full rank $r$ (see Thm. 1.3 and Remark 1.2). Consider SVD
$M=U\cdot\begin{bmatrix}\Sigma_{r}&0\\ 0&\Sigma_{\perp}\end{bmatrix}\cdot\begin{bmatrix}Q^{*}\\ Q_{\perp}^{*}\end{bmatrix}$,
write $C_{1}:=Q^{*}H$, apply Cor. 8.1, and deduce that
[…]
(i) Recall from Cor. 9.1 that
$\operatorname{Probability}\big{\{}\sigma_{r}(C_{1})=\sigma_{r}(Q^{*}H)\geq\frac{1}{4}\sqrt{l/n}\big{\}}>1-\exp\Big{(}-\frac{n}{72}\Big{)}-\exp\Big{(}-\frac{l-r}{20}\Big{)}.$ (9.2)
(ii) Let $H=U_{H}\Sigma_{H}V_{H}^{*}$ be a compact SVD. Then $\sigma_{r}(C_{1})\geq\sigma_{r}(Q^{*}U_{H})\sigma_{r}(H)$.
Similarly to (9.2) obtain that
$\operatorname{Probability}\big{\{}\sigma_{r}(C_{1})\geq\frac{1}{4}\sqrt{l/n}\cdot\sigma_{r}(H)\big{\}}$ $\geq\operatorname{Probability}\big{\{}\sigma_{r}(Q^{*}U_{H})\geq\frac{1}{4}\sqrt{l/n}\big{\}}$
$>1-\exp\Big{(}-\frac{n}{72}\Big{)}-\exp\Big{(}-\frac{l-r}{20}\Big{)}.$
Combine Thm. 8.1, equation $Q_{\perp}||_{2}=1$, and the bound $\sigma_{r}(C_{1})\geq\frac{1}{4}\sqrt{l/n}\cdot\sigma_{r}(H)$ and obtain
[…]
Bound the output errors of Algs. 7.3 and 3.4 by combining the estimates of this section and Sec. 8.2 and by transposing the input matrix $M$.
9.3 Output errors of Range Finder for a perturbed factor-Gaussian input
Assumption 9.2.
For an $r\times n$ Gaussian matrix $G$ and a constant matrix $A\in\mathbb{R}^{m\times r}$ of full rank $r<\min(m,n)$ define the matrices $B:=\frac{1}{\sqrt{n}}\cdot G$ and $\tilde{M}:=AB$ and call $M:=\tilde{M}+E$ a perturbed right factor-Gaussian matrix if the Frobenius norm of a perturbation matrix $E$ is sufficiently small in comparison to $\sigma_{r}(A)$, as we specify in the next theorem.
Theorem 9.2.
[Errors of Range Finder for a perturbed right factor-Gaussian matrix.] Suppose that we are given an $r\times n$ Gaussian matrix $G$ and constant matrices $H\in\mathbb{R}^{n\times l}$, $A\in\mathbb{R}^{m\times r}$, and $E\in\mathbb{R}^{m\times n}$ for $r\leq l<\min(m,n)$. Let $\tilde{M}$ be a right factor-Gaussian matrix of Assumption 9.2 and let $M=\tilde{M}+E$ for a perturbation matrix $E$. Apply Alg. 7.1 to the matrix $M$ with a test matrix $H$, and let $X$ and $Y$ denote the output matrices. Assume that $n>36r$ and l>22(r1)l>22(r-1).
(i) Suppose that the matrix $H$ has orthonormal columns and that $E||_{F}\leq\frac{\sigma_{r}(A)}{48\sqrt{n/l}+6}$. Then
$|M-XY|||/\tilde{\sigma}_{r+1}(M)\leq\sqrt{1+100n/l}$
with a probability no less than $1-\exp(-\frac{l-r}{20})-\exp(-\frac{n-r}{20})-\exp(-\frac{n}{72})$.
[…]
$|M-XY|||/\tilde{\sigma}_{r+1}(M)\leq\sqrt{1+100\kappa^{2}(H)n/l}$
with a probability no less than $1-\exp(-\frac{l-r}{20})-\exp(-\frac{n-r}{20})-\exp(-\frac{n}{72}).$
Proof.
Recall that the matrices $B$, ABAB, and $BH$ have full rank with probability 1 (see Thm. 1.3) and assume that they do have full rank (see Remark 1.2). Let
$M=\begin{bmatrix}U_{r}&U_{\perp}\end{bmatrix}\begin{bmatrix}\Sigma_{r}&0\\ 0&\Sigma_{\perp}\end{bmatrix}\begin{bmatrix}V_{r}^{T}\\ V_{\perp}^{T}\end{bmatrix}\textrm{ and }\tilde{M}=\begin{bmatrix}\tilde{U}_{r}&\tilde{U}_{\perp}\end{bmatrix}\begin{bmatrix}\tilde{\Sigma}_{r}&0\\ 0&0\end{bmatrix}\begin{bmatrix}\tilde{V}_{r}^{T}\\ \tilde{V}_{\perp}^{T}\end{bmatrix}$
be SVDs, where $V_{r}$ and $\tilde{V}_{r}$ are the matrices of the $r$ top right singular vectors of $M$ and $\tilde{M}$, respectively. Define $C_{1}=V_{r}^{T}H$ and $\tilde{C}_{1}:=\tilde{V}_{r}^{T}H$ as in (8.1). Now Thm. 8.1, that is, [30, Thm. 9.1], implies that
[…]
Recall that the columns of $V_{r}$ span the row space of an $r\times n$ Gaussian matrix $G$ and deduce from Corollary 9.1 that $\sigma_{r}(\tilde{C}_{1})>\frac{1}{4}\sqrt{l/n}$ with a probability no less than $1-\exp(-\frac{l-r}{20})-\exp(-n/72)$. It remains to estimate how the perturbation $E$ alters the leading right singular space of $\tilde{M}$ and how close is $\sigma_{r}(C_{1})$ to $\sigma_{r}(\tilde{C}_{1})$.
[…]
Claim (ii) of Thm. C.2 implies that whp the $r$-th singular value of $\tilde{M}=AB$ is not much less than the $r$-th singular value of $A$. Readily deduce from Cor. 9.1 that
$\operatorname{Probability}\{\sigma_{r}(\tilde{M})<\sigma_{r}(A)/3\}\leq\operatorname{Probability}\{\sigma_{r}(B)=\frac{1}{\sqrt{n}}\sigma_{r}(G)<1/3\}\leq e^{-(n-r)/20}.$
Hence $E||_{F}\leq\frac{\sigma_{r}(A)}{48\sqrt{n/l}+6}\leq\frac{\sigma_{r}(\tilde{M})}{16\sqrt{n/l}+2}$ with a probability no less than $1-e^{-(n-r)/20}$, and so $P||_{2}\leq\frac{1}{8}\sqrt{l/n}$ for some matrix $P$ of Lemma 1.2 such that $P||_{2}\leq\frac{1}{8}\sqrt{l/n}$.
Now let this holds, let $\sigma_{r}(\tilde{C}_{1})>\frac{1}{4}\sqrt{l/n}$, and deduce that
$\sigma_{r}(V_{r}^{T}H)$ $=\sigma_{r}\big{(}(I_{r}+P^{T}P)^{-1/2}(\tilde{V}_{r}^{T}+P^{T}\tilde{V}_{\perp}^{T})H\big{)}$ (9.3)
$\geq\sigma_{r}\big{(}(I_{r}+P^{T}P)^{-1/2}\big{)}\sigma_{r}(\tilde{V}_{r}^{T}H+P^{T}\tilde{V}_{\perp}^{T}H)$ (9.4)
$\geq\frac{1}{\sqrt{1+(\sigma_{1}(P))^{2}}}\big{(}\sigma_{r}(\tilde{C}_{1})-\sigma_{1}(P)\big{)}>\frac{1}{10}\sqrt{l/n}.$ (9.5)
Equality (9.3) holds because the matrix $\tilde{V}_{r}+\tilde{V}_{\perp}P$ is normalized by $(I_{r}+P^{T}P)^{-1/2}$ (see Remark 1.1) and has the same column span as $V_{r}$. By applying the union bound deduce that inequality (9.5) holds with a probability no less than $1-\exp(-\frac{l-r}{20})-\exp(-\frac{n-r}{20})-\exp(-\frac{n}{72})$.
To prove claim (ii), we essentially need to show that $\sigma_{r}(C_{1})=\sigma_{r}(V_{r}^{T}H)\geq\frac{1}{10}\sqrt{l/n}\cdot\sigma_{l}(H)$, and then the claim will follow readily from inequality (9.3). Let $H=U_{H}\Sigma_{H}V_{H}^{T}$ be compact SVD, such that $U_{H}\in\mathbb{R}^{n\times l}$ , $\Sigma_{H}\in\mathbb{R}^{l\times l}$, and $V_{H}\in\mathbb{R}^{l\times l}$, and obtain that $\sigma_{r}(\tilde{C}_{1})\geq\sigma_{r}(\tilde{V}_{r}^{T}U_{H})\sigma_{l}(\Sigma_{H})$ and
$\operatorname{Probability}\{\sigma_{r}(\tilde{C}_{1})<\frac{1}{4}\sqrt{l/n}\cdot\sigma_{l}(H)\}<\exp(-\frac{l-r}{20})+\exp(-\frac{n}{72}).$
Next bound $\sigma_{r}(C_{1})$ by showing that the column spaces of $V_{r}$ and $\tilde{V}_{r}$ are sufficiently close to one another if the perturbation $V_{r}-\tilde{V}_{r}$ is sufficiently small. Namely, assume that $E||_{F}\leq\frac{\sigma_{r}(A)}{12}$, and then the assumptions of Lemma 1.2 holds whp. By applying the same argument as in the proof of claim (i), deduce that
[…]
with a probability no less than $1-e^{-(n-r)/20}$. It follows that $P||_{2}\leq\frac{1}{8}\sqrt{l/n}\cdot\sigma_{l}(H)$ for some matrix $P$ of Lemma 1.2. Hence $\sigma_{r}(C_{1})\geq\frac{1}{10}\sqrt{l/n}\cdot\sigma_{l}(H)$ whp. ∎
10 Numerical tests
Next we cover our tests of dual superfast variants of Algs. 7.1 and 7.3 specified in Remark 7.2. As in Sec. 6 we applied the standard normal distribution function randn of MATLAB to generate Gaussian matrices. We apply MATLAB function ”svd()” to calculate the $\epsilon$-rank for $\epsilon=10^{-6}$ and perform the tests for Tables 10.110.3 on a 64-bit Windows machine with an Intel i5 dual-core 1.70 GHz processor by using custom programmed software in $C^{++}$ and compiled with LAPACK version 3.6.0 libraries, as in Sec. 6.
10.1 Input matrices
We generated the following classes of input matrices $M$ for testing LRA algorithms.
Class I: $M=U_{M}\Sigma_{M}V_{M}^{*}$, where $U_{M}$ and $V_{M}$ are the Q factors of the thin QR orthogonalization of $n\times n$ Gaussian matrices, $\Sigma_{M}=\operatorname{diag}(\sigma_{j})_{j=1}^{n}$; $\sigma_{j}=1/j,~j=1,\dots,r$, $\sigma_{j}=10^{-10},~j=r+1,\dots,n$ (cf. [H02, Sec. 28.3]), and $n=256,512,1024$. (Hence $M||_{2}=1$ and $M^{+}||_{2}=10^{10}$.)
Class II: (i) The matrices $M$ of the discretized single-layer Laplacian operator of [30, Sec. 7.1]: $S\sigma](x)=c\int_{\Gamma_{1}}\log{|x-y|}\sigma(y)dy,x\in\Gamma_{2}$, for two circles $\Gamma_{1}=C(0,1)$ and $\Gamma_{2}=C(0,2)$ on the complex plane. We arrived at the matrices $M=(m_{ij})_{i,j=1}^{n}$, $m_{i,j}=c\int_{\Gamma_{1,j}}\log|2\omega^{i}-y|dy$ for a constant $c$, M=1||M||=1 and the arc $\Gamma_{1,j}$ of $\Gamma_{1}$ defined by the angles in the range $\frac{2j\pi}{n},\frac{2(j+1)\pi}{n}]$.
(ii) The matrices that approximate the inverse of a large sparse matrix obtained from a finite-difference operator of [30, Sec. 7.2].
Class III: The dense matrices of five families from Class II of Sec. 6.2.
We used $1024\times 1024$ SVD-generated input matrices of class I having numerical rank $r=32$, $400\times 400$ Laplacian input matrices of class II(i) having numerical rank $r=36$, $408\times 800$ matrices having numerical rank $r=145$ and representing finite-difference inputs of class II(ii), and $1000\times 1000$ matrices of class III, having numerical rank 4, 6, 10, 12, and 25 for the matrices of the five families wing, baart, foxgood, shaw, and gravity, respectively.
10.2 Five families of Ultrasparse multipliers $H$
We generated our $n\times(r+p)$ sketch matrices $H$ for random $p=1,2,\dots,21$ by using 3-ASPH, 3-APH (see the end of Appendix B.2), and Random permutation matrices. When the overestimation parameter $p$ was considerable, we actually computed LRA of numerical rank larger than $r$, and so LRA was frequently closer to an input matrix than the optimal rank-$r$ approximation. Accordingly, the output error norms in our tests ranged from about $10^{-4}$ to $10^{4}$ relative to the optimal errors.
We obtained every 3-APH and every 3-ASPH matrix by applying three Hadamard’s recursive steps (B.1) followed by random column permutation defined by random permutation of the integers from 1 to $n$ inclusive. While generating a 3-ASPH matrix we also applied random scaling with a diagonal matrix $D=\operatorname{diag}(d_{i})_{i=1}^{n}$ where we have chosen the values of independent identically distributed (iid) random variables $d_{i}$ sampled under the uniform probability distribution from the set $\{-4,-3,-2,-1,0,1,2,3,4\}$.
We used the following families of sketch matrices $H$: (0) Gaussian (for control), (1) sum of a 3-ASPH and a permutation matrix, (2) sum of a 3-ASPH and two permutation matrices, (3) sum of a 3-ASPH and three permutation matrices, (4) sum of a 3-APH and three permutation matrices, and (5) sum of a 3-APH and two permutation matrices.
10.3 Test results
Tables 10.110.6 display the average relative error norm $\frac{\|M-\tilde{M}\|_{2}}{\|M-M_{nrank}\|_{2}}$ in our tests repeated 100 times for each class of input matrices and each size of an input matrix and multiplier $H$ for Alg. 7.1 or for each size of an input matrix and a pair of left-hand and right-hand multipliers $F$ and $H$ for Alg. 7.3.
In all our tests we applied the multipliers of the six families of the previous subsection.
Tables 10.110.3 display the average relative error norm for the output of Alg. 7.1; in our tests it ranged from about $10^{-3}$ to $10^{1}$. The numbers in parentheses in the first line of Tables 10.2 and 10.3 show the numerical rank of input matrices.
Tables 10.410.6 display the average relative error norm for the output of Alg. 7.3 applied to the same input matrices from classes I–III as in our experiments for Alg. 7.1.
In these tests we used $n\times\ell$ and $\ell\times m$ multipliers for $\ell=r+p$ and $k=c\ell$ for $c=1,2,3$ and random $p=1,2,\dots,21$.
SVD-generated Matrices Laplacian Matrices Finite Difference Matrices
Family No. Mean Std Mean Std Mean Std
Family 0 4.52e+01 5.94e+01 6.81e-01 1.23e+$0$ 2.23e+$0$ 2.87e+$0$
Family 1 3.72e+01 4.59e+01 1.33e+$0$ 2.04e+$0$ 8.22e+$0$ 1.10e+01
Family 2 5.33e+01 6.83e+01 1.02e+$0$ 2.02e+$0$ 4.92e+$0$ 4.76e+$0$
Family 3 4.82e+01 4.36e+01 7.56e-01 1.47e+$0$ 4.82e+$0$ 5.73e+$0$
Family 4 4.68e+01 6.65e+01 7.85e-01 1.31e+$0$ 3.53e+$0$ 3.68e+$0$
Family 5 5.45e+01 6.23e+01 1.03e+$0$ 1.78e+$0$ 2.58e+$0$ 3.73e+$0$
Table 10.1: Relative error norms in tests for matrices of classes I and II
wing (4) baart (6)
Family No. Mean Std Mean Std
Family 0 1.07e-03 6.58e-03 2.17e-02 1.61e-01
Family 1 3.54e-03 1.39e-02 1.37e-02 6.97e-02
Family 2 4.74e-03 2.66e-02 1.99e-02 8.47e-02
Family 3 1.07e-03 5.69e-03 1.85e-02 8.74e-02
Family 4 4.29e-03 1.78e-02 8.58e-03 5.61e-02
Family 5 1.71e-03 1.23e-02 3.66e-03 2.38e-02
Table 10.2: Relative error norms for input matrices of class III (of San Jose University database)
foxgood (10) shaw (12) gravity (25)
Family No. Mean Std Mean Std Mean Std
Family 0 1.78e-01 4.43e-01 4.07e-02 1.84e-01 5.26e-01 1.24e+$0$
Family 1 1.63e+$0$ 3.43e+$0$ 8.68e-02 3.95e-01 3.00e-01 7.64e-01
Family 2 1.97e+$0$ 4.15e+$0$ 7.91e-02 4.24e-01 1.90e-01 5.25e-01
Family 3 1.10e+$0$ 2.25e+$0$ 4.50e-02 2.21e-01 3.63e-01 1.15e+$0$
Family 4 1.23e+$0$ 2.11e+$0$ 1.21e-01 5.44e-01 2.36e-01 5.65e-01
Family 5 1.08e+$0$ 2.32e+$0$ 1.31e-01 5.42e-01 2.66e-01 8.22e-01
Table 10.3: Relative error norms for input matrices of class III (of San Jose University database)
SVD-generated Matrices Laplacian Matrices Finite Difference Matrices
$k$ Class No. Mean Std Mean Std Mean Std
$\ell$ Family 0 2.43e+03 1.19e+04 1.28e+01 2.75e+01 9.67e+01 1.48e+02
Family 1 1.45e+04 9.00e+04 8.52e+03 8.48e+04 7.26e+03 2.47e+04
Family 2 4.66e+03 2.33e+04 3.08e+01 4.07e+01 3.80e+02 1.16e+03
Family 3 2.82e+03 9.47e+03 2.42e+01 3.21e+01 1.90e+02 3.90e+02
Family 4 3.15e+03 7.34e+03 2.71e+01 4.69e+01 1.83e+02 2.92e+02
Family 5 2.40e+03 6.76e+03 2.01e+01 3.56e+01 2.31e+02 5.33e+02
$2\ell$ Family 0 5.87e+01 5.59e+01 7.51e-01 1.33e+$0$ 3.17e+$0$ 3.89e+$0$
Family 1 7.91e+01 9.86e+01 3.57e+$0$ 7.07e+$0$ 1.55e+01 2.39e+01
Family 2 5.63e+01 3.93e+01 3.14e+$0$ 4.50e+$0$ 5.25e+$0$ 5.93e+$0$
Family 3 7.58e+01 8.58e+01 2.84e+$0$ 3.95e+$0$ 4.91e+$0$ 6.03e+$0$
Family 4 6.24e+01 4.54e+01 1.99e+$0$ 2.93e+$0$ 3.64e+$0$ 4.49e+$0$
Family 5 6.41e+01 6.12e+01 2.65e+$0$ 3.13e+$0$ 3.72e+$0$ 4.54e+$0$
$3\ell$ Family 0 9.29e+01 3.29e+02 8.33e-01 1.54e+$0$
Family 1 5.58e+01 4.20e+01 3.09e+$0$ 4.08e+$0$
Family 2 5.11e+01 4.94e+01 1.70e+$0$ 2.08e+$0$
Family 3 6.70e+01 8.27e+01 2.35e+$0$ 2.96e+$0$
Family 4 5.36e+01 5.74e+01 2.14e+$0$ 3.76e+$0$
Family 5 4.79e+01 4.58e+01 1.81e+$0$ 2.94e+$0$
Table 10.4: Relative error norms in tests for matrices of classes I and II
wing (4) baart (6)
$k$ Class No. Mean Std Mean Std
$\ell$ Family 0 1.70e-03 9.77e-03 4.55e+$0$ 4.47e+01
Family 1 3.58e+02 3.58e+03 1.42e-01 9.20e-01
Family 2 2.16e-01 2.10e+$0$ 1.10e-02 6.03e-02
Family 3 7.98e-04 7.22e-03 4.14e-03 3.41e-02
Family 4 5.29e-03 3.57e-02 2.22e+01 2.21e+02
Family 5 6.11e-02 5.65e-01 3.33e-02 1.30e-01
$2\ell$ Family 0 7.49e-04 5.09e-03 5.34e-02 2.19e-01
Family 1 4.74e-03 2.32e-02 2.14e-02 1.33e-01
Family 2 3.01e-02 2.34e-01 1.26e-01 7.86e-01
Family 3 2.25e-03 1.38e-02 5.91e-03 2.63e-02
Family 4 3.94e-03 2.54e-02 1.57e-02 6.71e-02
Family 5 2.95e-03 1.47e-02 1.58e-02 1.20e-01
$3\ell$ Family 0 4.59e-03 2.35e-02 1.50e-02 7.09e-02
Family 1 5.96e-03 2.82e-02 7.57e-03 4.84e-02
Family 2 1.74e-02 1.06e-01 6.69e-03 2.97e-02
Family 3 3.07e-03 3.07e-02 1.16e-02 5.16e-02
Family 4 2.57e-03 1.47e-02 2.35e-02 9.70e-02
Family 5 4.32e-03 2.70e-02 1.36e-02 5.73e-02
Table 10.5: Relative error norms for input matrices of class III (of San Jose University database)
foxgood (10) shaw (12) gravity (25)
$k$ Class No. Mean Std Mean Std Mean Std
$\ell$ Family 0 5.46e+$0$ 1.95e+01 8.20e-01 4.83e+$0$ 8.56e+$0$ 3.33e+01
Family 1 8.51e+03 1.88e+04 1.12e+$0$ 5.75e+$0$ 1.97e+01 1.00e+02
Family 2 5.35e+03 1.58e+04 1.93e-01 1.51e+$0$ 8.79e+$0$ 4.96e+01
Family 3 6.14e+03 1.74e+04 4.00e-01 1.90e+$0$ 7.07e+$0$ 2.45e+01
Family 4 1.15e+04 2.33e+04 2.95e-01 1.71e+$0$ 4.31e+01 3.80e+02
Family 5 7.11e+03 1.87e+04 2.18e-01 9.61e-01 6.34e+$0$ 2.59e+01
$2\ell$ Family 0 2.70e-01 7.03e-01 5.54e-02 2.62e-01 5.34e-01 1.59e+$0$
Family 1 5.24e+02 5.19e+03 4.67e-02 2.35e-01 1.38e+01 1.31e+02
Family 2 2.45e+$0$ 3.47e+$0$ 8.31e-02 6.37e-01 5.47e-01 1.69e+$0$
Family 3 2.43e+$0$ 3.74e+$0$ 1.24e-01 8.52e-01 5.10e-01 1.24e+$0$
Family 4 2.17e+$0$ 2.92e+$0$ 1.76e-01 8.76e-01 2.60e-01 7.38e-01
Family 5 2.10e+$0$ 3.34e+$0$ 1.26e-01 5.99e-01 5.68e-01 1.46e+$0$
$3\ell$ Family 0 2.62e-01 8.16e-01 4.49e-02 1.64e-01 4.59e-01 1.38e+$0$
Family 1 2.72e+$0$ 4.60e+$0$ 6.84e-02 3.43e-01 3.44e-01 8.60e-01
Family 2 2.42e+$0$ 3.92e+$0$ 8.26e-02 5.38e-01 6.89e-01 2.15e+$0$
Family 3 3.22e+02 3.20e+03 6.06e-02 2.95e-01 5.26e-01 1.17e+$0$
Family 4 1.91e+$0$ 3.36e+$0$ 6.61e-02 3.36e-01 6.19e-01 1.54e+$0$
Family 5 2.73e+$0$ 6.90e+$0$ 5.72e-02 2.39e-01 7.22e-01 1.59e+$0$
Table 10.6: Relative error norms for input matrices of class III (of San Jose University database)
11 Conclusions
Next we list some promising directions for further study of superfast LRA.
[44] proposed and tested iterative refinement of LRA, and [21] has greatly advanced that study incorporating the superfast algorithms that we studied here.
[20] proposed, analyzed, and tested superfast algorithms for matrix norm estimation.
Least squares approximation of multivariate functions (LSAMF) was reduced in [1] to the solution of Linear Least Squares Problem (LLSP) and further to LRA. Namely, [1] reduced LSAMF to LLSP, which we state as follows: for given $M\in\mathbb{C}^{m\times n}$ and ${\bf b}\in\mathbb{C}^{m\times 1}$ where $m>n$, compute or approximate a vector
[…]
Triangular inequality implies that for an LRA $\tilde{M}$ of $M$ we have
[…]
[1] showed that $(\tilde{M}-M){\bf y}||=0$ in the case of LSAMF problem, and so we can solve it superfast if LRA can be computed superfast, e.g., with our algorithms.
Indeed, suppose that our superfast LRA algorithm has computed $P$ and $S$ such that $\tilde{M}=PS\in\mathbb{R}^{m\times n}$, $P\in\mathbb{R}^{m\times r}$, $S\in\mathbb{R}^{r\times n}$ and $r\ll\min\{m,n\}$. Then compute a solution
[…]
Appendix
Appendix A Small families of hard inputs for superfast LRA
Any algorithm not involving all entries of an input matrix fails on the following input families.
Example A.1.
Let $\Delta_{i,j}$ denote an $m\times n$ matrix of rank 1 filled with 0s except for its $(i,j)$th entry filled with 1. The $mn$ such matrices $\{\Delta_{i,j}\}_{i,j=1}^{m,n}$ form a family of $\delta$-matrices. We also include the $m\times n$ null matrix $O_{m,n}$ filled with 0s into this family. Now fix any superfast algorithm; it does not access the $(i,j)$th entry of its input matrices for some pair of $i$ and $j$. Therefore, it outputs the same approximation of the matrices $\Delta_{i,j}$ and $O_{m,n}$, with an undetected error at least 1/2. Arrive at the same conclusion by applying the same argument to the set of $mn+1$ small-norm perturbations of the matrices of the above family and to the $mn+1$ sums of the latter matrices with any fixed $m\times n$ matrix of low rank. Finally, the same argument shows that a posteriori estimation of the output errors of an LRA algorithm applied to these input families cannot run superfast.
The example covers randomized LRA algorithms as well. Indeed, let an LRA algorithm miss an entry of an input matrix with a probability $p>0$. Apply this algorithm to two matrices of low rank whose difference at this entry is equal to a large constant $C$. Then with a probability $p$ the algorithm has error at least $C/2$ at this entry for a least one of these two matrices.
Appendix B LRA with dense and Abridged SRHT random multipliers
B.1 The generalized Nyström LRA algorithm (briefly)
The generalized Nyström algorithms of [60, 39] specify Alg. 7.3 for random sketching LRA.131313[39] modifies the algorithms of [60] to improve numerical stability of rank-$r$ approximation of a matrix $M\in\mathbb{C}^{m\times n}$ for $r\leq n\leq m$, but in our tests we departed from [60] and accelerated it by using Abridging SRHT multipliers $F$ and $H$. In the best studied case the multipliers $F$ and $H$ are Gaussian random matrices, filled with independent standard Gaussian (normal) random variables; then for $k=2r+1$ and $\ell=2k$ the expected Frobenius error norm of the error matrix MXYZM-XYZ is within a factor of 2 from optimal (not exceeding the optimal one for rank-$\frac{k}{2}$ approximation). The algorithm involves $O((k+\ell)mn)$ flops, $mn$ entries of $M$, and about $km+n\ell+k\ell$ other scalars, overwriting some entries of $M$.
With SRHT or SRFT (rather than Gaussian) multipliers $F$ and $H$, one obtains LRA (with a little higher error probability) by using $O(mn\log(k\ell)))$ flops, $mn$ entries of $M$, and $mk\log(k)+n\ell\log(\ell)$ other scalars, also overwriting some entries of $M$.
Like the other random sketching LRA algorithms, this one is superfast except for the stage of the computation of the sketches $FM$ and $MH$ and can be readily modified to bound the relative error of the output LRA by any $\epsilon>0$, with estimated sketches’ size increasing fast as $\epsilon\mapsto 0$.
B.2 Abridged SRHT matrices
With sparse subspace embedding [10, 12], [61, Sec. 3.3], [37, Sec. 9] one obtains significant acceleration but still does not yield superfast algorithms. According to [32], such acceleration tends to make the accuracy of the output LRAs somewhat less reliable, although [12] partly overcomes this problem for incoherent matrices. One can multiply a matrix by Subsampled Randomized Hadamard or Fourier Transform (SRHT or SRFT) dense matrices towards incoherence [12], but this step is not superfast. We make it superfast by applying Abridged SRHT multipliers, specified below. Thm. 9.1 provides semi-empirical support for the output accuracy of the resulting LRA. [48] covers earlier formal and empirical study of Abridged SRHT matrices and their application to LRA.
We proceed by means of abridging the classical recursive processes of the generation of SRHT matrices, obtained from the $n\times n$ dense matrices $H_{n}$ of Walsh-Hadamard transform for $n=2^{t}$ (cf. [35, Sec. 3.1]). Namely, we write
$H_{d,0}:=I_{n/2^{d}},~H_{d,i+1}:=\begin{pmatrix}H_{d,i}&H_{d,i}\\ H_{d,i}&-H_{d,i}\end{pmatrix}~{\rm for}~i=0,1,\dots,d-1.$ (B.1)
The $n\times n$ matrix $H_{d,d}$ is orthonormal up to scaling and has $2^{d}$ nonzero entries in every row and column. Now define $l\times n$ matrix $R$ of uniform random sampling of $l$ out of $n$ columns, the $n\times n$ diagonal matrix $D$ whose $n$ diagonal entries are independent random signs, i.e., random variables uniformly distributed on the pair $\{\pm 1\}$, and the $d$-Abridged subsampled randomized Hadamard transform (SRHT) matrix $\sqrt{2^{d}/l~}RH_{d,d}D$, with $2^{d}$ nonzero entries $\pm\sqrt{2^{d}/l}$ per column. We can multiply $M\in\mathbb{C}^{m\times n}$ by that matrix by involving $m2^{d}$ entries of $M$ and $(2^{d}-1)m$ flops.
Likewise, by sampling $k$ random rows of $H_{d,d}$ and multiplying the resulting matrix by $\pm\sqrt{2^{d}/k}D$ obtain $d$-Abridged $k\times m$ SRHT matrix having $2^{d}$ nonzero entries $\pm\sqrt{2^{d}/k}$ per row. We can multiply it by $M$ by involving $n2^{d}$ entries of $M$ and $(2^{d}-1)n$ flops, where a flop amounts to addition, subtraction, or multiplications by $\sqrt{2^{d}/l}$ or $\sqrt{2^{d}/k}$. In particular this means $8n$ (resp. $8m$) entries of $M$ and $7n$ (resp. $7m$) flops in our numerical tests, where we always fix $d=3$.
A $t$-Abridged SRHT matrix turns into an SRHT matrix.
Table B.1 displays the computational cost of randomized rank-$r$ approximation by means of the algorithm of [39], when it uses Gaussian, SRHT and $d$-Abridged SRHT multipliers $F$ and $H$.
Multipliers $F$ and $H$ Input entries involved Other entries involved Flops involved
Gaussian mnmn $km+n\ell+kl$ $O((k+l)mn)$
SRHT mnmn $O(mk\log(k)+nl\log(l))$ $O(mn\log(kl))$
$d$-Abridged SRHT $2^{d}(m+n)$ $2^{d}(k+l)$ $(2^{d}-1)(m+n)$
Table B.1: Computational cost of the generalized Nyström algorithm of [39] with Gaussian, SRHT, and $d$-Abridged SRHT multipliers $F$ and $H$.
We also refer to $d$-abridged SRHT matrices as $d$-Abridged Scaled and Permuted Hadamard $d$-ASPH matrices and call them $d$-APH matrices if random scaling is omitted, that is, if the multiplier $D$, filled with $\pm 1$, is replaced with the identity matrix, and then randomization only comes from random permutation matrix.
Appendix C The spectral norms of a Gaussian matrix and its pseudo inverse
Hereafter $\Gamma(x)=\int_{0}^{\infty}\exp(-t)t^{x-1}dt$ denotes the Gamma function; $\nu_{p,q}$ and $\nu^{+}_{p,q}$ denote the random variables representing the spectral norms of a $p\times q$ Gaussian random matrix and its Moore-Penrose pseudo inverse, respectively.
Theorem C.1.
[Spectral norms of a Gaussian matrix. See [17, Thm. II.7].]
[…]
Theorem C.2.
[Spectral norms of the pseudo inverse of a Gaussian matrix.]
(i) Probability $\{\nu_{m,n}^{+}\geq m/x^{2}\}<\frac{x^{m-n+1}}{\Gamma(m-n+2)}$ for $m\geq n\geq 2$ and all positive $x$,
(ii) Probability $\{\nu_{m,n}^{+}\geq t\frac{e\sqrt{m}}{m-n+1}\}\leq t^{n-m}$ for all $t\geq 1$ provided that $m\geq 4$,
(iii) $\mathbb{E}(\nu^{+}_{m,n})\leq\frac{e\sqrt{m}}{m-n}$ provided that $m\geq n+2\geq 4$,
Proof.
See [8, Proof of Lemma 4.1] for claim (i), [30, Prop. 10.4 and Eqns. (10.3) and (10.4)] for claims (ii) and (iii), and [57, Thm. 3.3] for claim (iv). ∎
Thm. C.2 implies reasonable probabilistic upper bounds on the norm $\nu_{m,n}^{+}$, even where the integer |mn||m-n| is close to 0; whp the upper bounds of Thm. C.2 on the norm $\nu^{+}_{m,n}$ decrease very fast as the difference |mn||m-n| grows from 1.
Appendix D Randomized pre-processing of LRA
The following simple results from [47, Section 8.2] show that pre-processing with Gaussian multipliers $X$ and $Y$ transforms any matrix that admits $LRA$ into a perturbation of a factor-Gaussian matrix. We write $A\preceq B$ (resp. $A\succeq B$) to show that $A$ is statistically less (resp. greater) than or equal to $B$.
Theorem D.1.
Consider five integers $k$, $l$, $m$, $n$, and $r$ satisfying the bounds $r\leq k\leq m,~r\leq l\leq n$, an $m\times n$ well-conditioned matrix $M$ of rank $r$, $k\times m$ and $n\times l$ Gaussian matrices $G$ and $H$, respectively, and the norms $\nu_{p,q}$ and $\nu_{p,q}^{+}$ of Appendix C. Then
(i) $GM$ is a left factor-Gaussian matrix of expected rank $r$ such that
[…]
(ii) $MH$ is a right factor-Gaussian matrix of expected rank $r$ such that
[…]
(iii) $GMH$ is a two-sided factor-Gaussian matrix of expected rank $r$ such that
[…]
Remark D.1.
According to this theorem we can readily extend our results on $LRA$ of perturbed factor-Gaussian matrices to all matrices that admit $LRA$ and are pre-processed with Gaussian multipliers. We can perform such pre-processing superfast if we replace Gaussian multipliers with 3-abridged SRHT orthogonal multipliers; this replacement has little affected the output accuracy of LRA for a large class of matrices $M$ in the tests in [21].
Appendix E Proof of Thm. 2.1
Proof.
$\sigma_{j}(G)\leq\sigma_{j}(M)$ for all $j$ because $G$ is a submatrix of $M$. Hence $\epsilon$-$(G)\leq\epsilon-rank$$(M)$ for all nonnegative $\epsilon$, and in particular $\operatorname{rank}(G)\leq\operatorname{rank}(M)$.
Now let $M=M^{\prime}=CUR$. Then clearly
$\operatorname{rank}(M)\leq\operatorname{rank}(U)=\operatorname{rank}(G_{r}^{+})=\operatorname{rank}(G_{r})\leq\operatorname{rank}(G),$
Hence
$\operatorname{rank}(G)\geq\operatorname{rank}(M),~{\rm and~so}~\operatorname{rank}(G)=\operatorname{rank}(M)~{\rm if}~M^{\prime}=M.$
It remains to deduce that
$M=CG_{r}^{+}R~{\rm if}~\operatorname{rank}(G)=\operatorname{rank}(M):=r,$
but in this case $G_{r}=G$, and so
$\operatorname{rank}(CG_{r}^{+}R)=\operatorname{rank}(C)=\operatorname{rank}(R)=r.$
Hence the rank-$r$ matrices $M$ and $CG_{r}^{+}R$ share their rank-$r$ submatrices $C$ and $R$. ∎
Remark E.1.
Can we extend the theorem by proving that $M^{\prime}\approx M$ if and only if $\epsilon$-$(G)=\epsilon-rank$$(M)$ for a small positive $\epsilon$? The “only if” claim cannot be extended, e.g., for
$M=\begin{pmatrix}1&0&0\\ 0&\epsilon&0\\ 0&1&0\end{pmatrix}$,
$\epsilon\approx 0$, and the $2\times 2$ leading submatrix $G$ of $M$. Indeed, $\operatorname{rank}(M)=\operatorname{rank}(G)=2$, and so Thm. 2.1 implies that $M^{\prime}=M$, while $\epsilon$-$\operatorname{rank}(M)=2>\epsilon$-$\operatorname{rank}(G)=1.$
Acknowledgements: Our work has been supported by NSF Grants CCF–1116736, CCF–1563942 and CCF–1733834 and PSC CUNY Awards 66720-00 54 and 68452-00 56. We are also grateful to E. E. Tyrtyshnikov for the challenge of formally supporting empirical power of C–A iterations, to N. L. Zamarashkin for his comments on [46], and to S. A. Goreinov, I. V. Oseledets, A. Osinsky, E. E. Tyrtyshnikov, and N. L. Zamarashkin for reprints and pointers to relevant bibliography.